### v1

# Quenched coalescent for diploid population models with selfing and overlapping generations

Abstract
We introduce a general diploid population model with self-fertilization and possible overlapping generations, and study the genealogy of a sample of $n$ genes as the population size $N$ tends to infinity. Unlike traditional approach in coalescent theory which considers the unconditional (annealed) law of the gene genealogies averaged over the population pedigree, here we study the conditional (quenched) law of gene genealogies given the pedigree. We focus on the case of high selfing probability and obtain that this conditional law converges to a random probability measure, given by the random law of a system of coalescing random walks on an exchangeable fragmentation-coalescence process of Ber (04). This system contains the system of coalescing random walks on the ancestral recombination graph as a special case, and it sheds new light on the site-frequency spectrum (SFS) of genetic data by specifying how SFS depends on the pedigree. The convergence result is proved by means of a general characterization of weak convergence for random measures on the Skorokhod space with paths taking values in a locally compact Polish space.
1 Introduction
Coalescent processes have been widely used as models of gene genealogies that describe the ancestral structure of a sample of $n$ genes, when the total population size $N$ is sufficiently large. The Kingman coalescent Kin (82), for instance, has been enormously impactful in the study of natural genetic variation in populations Wak (09). Its power stems from its remarkable robustness; indeed, a large number of population models were shown to have the Kingman coalescent or its variant as their scaling limit as $N$ tends to infinity Möh (98). Other models such as the coalescent with asynchronous multiple mergers DK (99); Sag (99); Pit (99) and the coalescent with simultaneous multiple mergers MS (03); Sch (00); BLS (18) have also been discovered as scaling limits under exchangeable models, when the number of offspring per individuals has very high variance, and have been applied to a number of different species FKM+ (23). The simultaneous multiple-mergers coalescent, also called the $\Xi$-coalescent, is most relevant to this paper.
Traditionally, coalescent models are obtained by taking average over the population pedigree, the graph that represents the total history of reproductive relationships in the population. Namely, implicit in the approach of papers in classical coalescent theory Kin (82); Sag (99); Pit (99); Sch (00); BLS (18) is an annealing over all realizations of the pedigree to describe the distribution of gene genealogies across unlinked loci in the genome. However, this tradition of averaging over the pedigree is questionable because there is only one population pedigree, and all genetic information across loci is passed through this same pedigree. For example, in order for unlinked loci to have independent genealogies (as they must by definition), under this averaging they would also need to have independent pedigrees. This is problematic because even unlinked loci are subject to the same pedigree.
This conceptual flaw was not unrecognized BNA (90); WKLR (12); WKW (16); WBLW (17); Ral (19), but perhaps luckily, it turned out that the Kingman coalescent can still be applied under the standard assumptions of neutral coalescent model because then the conditional limit is still equal to the unconditional limit Tyu (15). However, when there are simultaneous multiple mergers, the conditional genealogy can be different from the unconditional genealogy DFBW (24).
We believe that, at least at the conceptual level, coalescent theory should start by conditioning on the pedigree. This has been the basis of the mathematically rigorous works Tyu (15); DFBW (24); NWF (25); ABFW (25) that collectively marked the emerging quenched-coalescent theory. We now give a brief account of these works before describing the main contribution of the present paper.
In ABFW (25), the authors study the diploid Cannings model introduced in BLS (18), where selfing is excluded. This is a model in which the offspring distribution, described by the matrix $(V_{ij})$ of the number of offspring between all pairs of parents, is invariant under permutations of the parents. Under the same condition on the offspring distribution that guaranteed annealed convergence to the $\Xi$-coalescent holds, the authors in ABFW (25) showed that the genealogies conditioned on the pedigree converge to an inhomogeneous $(\Psi,c_{\rm pair})$-coalescent, where $\Psi$ is a Poisson point process on $0,\infty)\times(\Delta\setminus\{\bf 0\})$ with intensity $dt\otimes\frac{1}{\langle x,x\rangle}\Xi(dx)$. This limiting process consists of the independent superposition of a Kingman coalescent with rate $c_{\rm pair}$ and multiple merger events whose the timing and intensity are specified by $\Psi$. The result in ABFW (25) generalizes earlier work in Tyu (15) and DFBW (24).
While the Cannings model in BLS (18); ABFW (25) is quite general and captures a wide range of reproductive variance, it has two limitations from a biological perspective. Firstly, it does not allow for selfing (self-fertilization), an important evolutionary force found in taxa including eukaryotic microbes and marine invertebrates (SR, 17; YSH, 23) and is common in plants (AG, 89; HBG, 17; TWL+, 23). The empirical distribution of selfing probabilities among plant species is markedly bimodal, with relatively fewer species having $s\in(0.2,0.8)$ and some species reaching as high as 0.99 SL (85); AG (89); VK (01); BF (05); GKE (05); SAAMT (20). Secondly, the model assumes non-overlapping generations, a simplification that excludes biologically realistic scenarios in which individuals from different age cohorts may reproduce simultaneously. Overlapping generations are common in many natural populations and can introduce temporal correlations and ancestral dependencies not captured by discrete-generation models.
The recent work NWF (25) introduces a diploid Moran model with a high selfing probability $\alpha_{N}$. Quenched coalescent limit for the coalescence time of a sample of size $n=2$ was obtained for that Moran model. Conditioning coalescence on the pedigree reveals three markedly different behaviors, depending on how quickly $\alpha_{N}\to 1$. The critical case, which we call ‘limited outcrossing’, is when $1-\alpha_{N}$ is of order $1/N$.
This paper aims to contribute to building the emerging quenched-coalescent theory, by going beyond the existing mathematical work Tyu (15); DFBW (24); NWF (25); ABFW (25). We introduce a general diploid model with overlapping generations that extends the haploid model of SW (08) and the diploid models of BLS (18) and NWF (25). Our model explicitly incorporates a selfing rate parameter and strictly contains the diploid Moran-type models with selfing in NWF (25); CLJ (22); Lin (09) as special cases. Our main result significantly strengthens the analysis in NWF (25), in the ‘limited outcrossing’ regime, by considering arbitrary sample size $n$ and a substantially broader class of models. Moreover, it complements the results of ABFW (25); Tyu (15); DFBW (24) by identifying a distinct limiting conditional coalescent process that arises in the high selfing regime.
The new class of coalescent models described herein is a family of coalescing random walks on a directed random graph, which we call a $Q$-$\lambda$ graph in this paper. This random graph is a subgraph of an exchangeable fragmentation-coalescence process (EFC) introduced in Ber (04), where $Q$ is the rate matrix for coalescence and $\lambda$ is the fragmentation rate at which each node spits into two. The parameter $\lambda$ is the relative rate of outcrossing to coalescence as $N\to\infty$. In the regime where the classical (annealed) limit is a time-rescaled Kingman, we find that these $Q$-$\lambda$ graphs correspond to the ancestral recombination graphs (also known as ancestral selection graph) Hud (83); GM (97); KN (97) which are used to describe the genealogy of a sample of genome sequences with recombination.
The proof of the main result involves a novel characterization of convergence of random measures on the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for any locally compact Polish space $E$, and a general theorem about quenched convergence of Markov processes on a suitably enriched space of partitions. These results are then combined with a combinatorial argument to characterize the scaling limits of our diploid Sargasyan-Wakeley model. These results also allow us to strengthen the results of (ABFW, 25, Theorem 3.8) from quenched convergence in finite-dimensional distribution to weak convergence in distribution of random measures.
From the application standpoint, the main motivation of coalescent theory (and our quenched-coalescent theory) is to describe the patterns of genetic variation expected under various biological scenarios to provide frameworks for statistical inference about past events and processes affecting populations. In a given population, gene genealogies are tree structures which emerge from tracing the ancestral lines of these sample backwards in time until the most recent common ancestor (MRCA). Mutations occurred in the past, along the ancestral lines, result in genetic diversity among the sample. For example, the site-frequency spectrum (SFS) is a commonly used measure of genetic variation upon which statistical inferences are based BWSH (01); Ach (09); EBBF (15); LF (15); GK (16); FLW+ (17); FKM+ (23). For a sample of $n$ genomes, the SFS records the number of polymorphic sites where a mutant base is found in $r\in\{1,\ldots,n-1\}$ copies Taj (89); BHK+ (95); Fu (95). Because per-site mutation rates are typically very small and the number of sites is large, the SFS is taken to reflect the total length of branches in the gene genealogy with $r$ descendants in the sample, or which are ancestral to $r$ of the samples. Different biological phenomena, such as population growth and natural selection, lead to different coalescent predictions for the SFS. Our main result sheds new light on the SFS of genetic data, by specifying how SFS depends on the pedigree. More precisely, under the conditional coalescent, the SFS should be viewed as a conditional SFS given the single pedigree as a latent variable. In Figure 5, we provide simulations for the conditional SFS given 5 different pedigrees and thus illustrate how the pedigree can impact the SFS of the data.
Organization of this paper. Section 2 introduces our diploid exchangeable model with self-fertilization and overlapping generations. An annealed version of our main result, akin to classical results of coalescent theory such as those of Kin (82); Möh (98); Pit (99); Sag (99); Sch (00), is presented in Section 3. Section 4 presents the main quenched convergence result for the coalescent of our exchangeable model model conditional on the pedigree. In Section 5 we present applications of our main convergence result and illustrate how the SFS can depend strongly on the pedigree. The remainder of the paper focuses on the tools and calculations necessary to prove the main results: weak convergence of random measures in distribution in Section 6, and a suitable notion of convergence of the random pedigree and continuity of the coalescent law’s dependence thereon in Section 7.
2 An exchangeable diploid model with selfing and overlapping generation
2.1 The model
In this paper, we consider a diploid, monoecious, panmictic (well-mixed and randomly mating) population of constant size $N$, evolving in discrete time-steps with overlapping generations. This model generalizes the diploid Cannings model in ABFW (25); BLS (18) to allow overlapping generations and selfing, and extends the haploid model in SW (08) to diploid populations while supporting general offspring distributions.
Precisely, the population model is specified by a deterministic number $\alpha_{N}\in[0,1]$ and the joint distribution of a random variable $K_{N}$ and a random symmetric matrix $\textbf{V}=(V_{i,j})_{1\leq i,j\leq N}$. The number $\alpha_{N}$ represents the selfing probability for each offspring, and $K_{N}$ and V represent, at an arbitrary timestep, the total offspring number and the pairwise offspring numbers respectively. We assume the following:
  • $K_{N}$ and all entries of V take values in $\{0,1,2,\ldots,N\}$.
  • The total number of offspring satisfies $\sum_{i\leq j}^{N}V_{i,j}=K_{N}$.
  • The full matrix V is exchangeable, i.e.,
    $\left(V_{i,j}\right)_{1\leq i,j\leq N}\overset{d}{=}\left(V_{\sigma(i),\sigma(j)}\right)_{1\leq i,j\leq N}$
    for any permutation $\sigma$ of $N]=\{1,2,\ldots,N\}$.
We consider discrete time-steps indexed by $k\in\mathbb{Z}_{+}=\{0,1,2,\ldots\}$, where $k=0$ is the present, $k=1$ the previous time-step, and so on backward into the past. Let $\big\{(K^{(k)}_{N},\textbf{V}^{(k)})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random variables that have the same distribution as $(K_{N},\textbf{V})$. Reproduction events in different time-steps are taken to be independent, and are described as follows:
For each $k\in\mathbb{Z}_{+}$, $K^{(k)}_{N}$ individuals are chosen uniformly without replacement, among the $N$ individuals from the $k$-th time-step, to be children. To describe parentage, we suppose that the individuals at each time-step are labeled by $[N]$, and we let $V^{(k)}_{i,j}$ be the number of offspring produced by the pair of individuals $(i,j)$ in the $(k+1)$-th time-step in the past. The $K_{N}^{(k)}$ individuals defined by the entries of $\textbf{V}^{(k)}$ constitute the new individuals in time-step $k$. Each individual has two sets of chromosomes, as it is a diploid population. Genetic lineages at an autosomal locus are transmitted according to Mendel’s law of random segregation, which means each gene copy in the offspring chooses independently from the two gene copies in the parent from which the gene copy is inherited.
The remaining $N-K_{N}^{(k)}$ individuals are carried over unchanged (without random segregation) from time-step $k+1$, which can induce overlapping generations.
We further assume that each of the $K^{(k)}_{N}$ children independently chooses to have either a single parent (selfing) or two distinct parents (outcrossing), with probabilities $\alpha_{N}$ and $1-\alpha_{N}$ respectively. Therefore, conditional on $K_{N}^{(k)}$, the number of selfed offspring $S_{N}^{(k)}:=\sum_{i=1}^{N}V_{i,i}^{(k)}$ is an independent binomial random variable: $\mathrm{Bin}(K_{N}^{(k)},\alpha_{N})$. In other words, we assume that $(K_{N},\textbf{V})$ satisfies
$\sum_{i=1}^{N}V_{i,i}\sim\mathrm{Bin}(K_{N},\alpha_{N})\qquad\text{given }K_{N}.$
Figure 1: An illustration of our population model between time-steps $k+1$ and $k$ in the past, with size $N=6$. The number of children, whose edges are colored in black, is $K_{N}=3$. Further, $S_{N}=2$ as two of the children have reproduced by selfing. Individuals $1,2,4$, whose edges are marked in orange, simply persisted between consecutive time-steps and are not children. To demonstrate how to read the $V_{i}$ and $V_{i,i}$ from the pedigree, we see that $V_{1}=0$ as the first individual has no offspring. $V_{2}=2$ as it has the third and fourth individuals as offspring. $V_{2,2}=1$ as the third individual is a child via offspring while the fourth is via selfing.
See Figure 1 for a realization of one time-step of this process when $N=6$.
The population model described above includes two overlapping but distinct classes of models in the literature as special cases.
Example 2.1 (The diploid Cannings model in BLS (18)).
When $K_{N}=N$ (non-overlapping generations) and $\alpha_{N}=0$ (no selfing), our model reduces to the Cannings model introduced in BLS (18) and studied in ABFW (25).
Example 2.2 (A diploid Sargasyan-Wakeley model).
Our model also generalizes the model in SW (08) in two ways: to incorporate more general offspring distribution, and from haploid to diploid.
Suppose the random symmetric matrices $\{V^{(k)}\}_{k\in\mathbb{Z}_{+}}$ satisfy the following extra assumption. Namely, their distribution is parametrized by also a random number $P_{N}\in\{2,3\ldots,N\}$ representing the number of potential parents at a timestep.
Let $\big\{(K^{(k)}_{N},\,P^{(k)}_{N})\big\}_{k\in\mathbb{Z}_{+}}$ be a sequence of i.i.d. random vectors that have the same distribution as $(K_{N},\,P_{N})$. For each $k\in\mathbb{Z}_{+}$, as before $K^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $k$-th time-step to be children. Then, suppose that (i) $P^{(k)}_{N}$ individuals are chosen uniformly without replacement from the $(k+1)$-th time-step to be potential parents of the $K^{(k)}_{N}$ children, and (ii) the actual parent(s) of the child is (are) chosen uniformly without replacement from the $P^{(k)}_{N}$ potential parent(s).
When $K_{N}=P_{N}=N$, this is the Wright-Fisher model with selfing considered in Möh (98); ND (97) and KDWF (23) with free recombination ($r_{N}=1/2$) and $\alpha_{N}=s_{N}$. The model in DFBW (24) with $\psi=1$ is exactly the case when $K_{N}=N$ deterministically and $P_{N}$ is a random variable taking values $2$ and $N$ with probabilities $\lambda/N^{\theta}$ and $1-\lambda/N^{\theta}$ respectively.
The model of NWF (25) corresponds to the special case when $K_{N}=1$ deterministically. When $\alpha_{N}=\alpha\in[0,1)$ for all $N$, this corresponds to the diploid Moran model with selfing considered in Lin (09). When $\alpha=0$ this corresponds to the model as in CLJ (22). The model in BBE (13) corresponds to our case when $\alpha_{N}=0$, and when $(K_{N},P_{N})$ is equal to $(1,2)$ and $([\psi N],2)$ with probabilities $1-\varepsilon_{N}$ and $\varepsilon_{N}$ respectively. There is a slight difference between these models and our model here, as the indices of the potential parents and those of the offspring are enforced to be disjoint in these models.
To simplify notation, we shall omit the subscript $N$ when there is no confusion. For example, we will write $(\alpha,K,S,K^{(k)})$ instead of $(\alpha_{N},K_{N},S_{N},K^{(k)}_{N})$.
Let $V_{i}:=\sum_{j=1}^{N}V_{i,j}$ is the total number of offspring for individual $i$, and $\tilde{V}_{i}:=V_{i,i}+V_{i}$. Then $\sum_{i=1}^{N}\tilde{V}_{i}=2K$ and the total genetic contribution of the $i$th individual to the next time step is
$\frac{\tilde{V}_{i}}{2N}:=\frac{V_{i,i}+V_{i}}{2N}=\frac{1}{2N}\Bigl(2V_{i,i}+\sum_{j\neq i}V_{i,j}\Bigr),$ (2.1)
because the genetic contribution of a parent by selfing is twice that via outcrossed offspring.
In the absence of selfing ($\alpha_{N}=0$), we have $\tilde{V}_{i}=V_{i}$, recovering the offspring-frequency weights of MS (03) (see also BLS (18); ABFW (25)). These frequencies provide a convenient device to express one-step coalescence probabilities (e.g. pairwise and triple collisions) as explicit polynomials in $\{\tilde{V}_{i}\}$ with Mendelian coefficients.
2.2 The pedigree as important latent variable of the population
The population dynamics described in Section 2.1 give rise to a random directed graph $\mathcal{G}_{N}$ that encodes the population pedigree, i.e. the reproductive relationships among all individuals. We give a formal definition below and offer the left panel of Figure 2 as an illustration.
Each individual is diploid and carries two copies of each autosomal locus. Genetic lineages are transmitted through the population pedigree via Mendelian segregation: each gene copy in the offspring randomly inherits one allele from the corresponding gene copies in the designated parent (or one of the two gene copies in the case of selfing). The pedigree is described explicitly in Definition 2.3.
Figure 2: (Population process on the right and the corresponding pedigree on the left). A realization of our population process with $N=5$ individuals from time-steps $k=0$ to $5$ in the past (right panel). The corresponding pedigree is shown on the left. Here we consider the Moran model in NWF (25), where $K_{N}=1$ deterministically. The black lines correspond to reproductive relationships while the yellow lines correspond to an individual persisting from one time-step to the next. The yellow lines give rise to overlapping generations.
Definition 2.3 (Pedigree).
The population pedigree, or simply the pedigree, refers to an undirected multi-graph $\mathcal{G}_{N}$ with vertex set $N]\times\mathbb{Z}_{+}$, where a vertex $(\ell,k)$ represents the individual with index $\ell$ at the $k$-th time-step. The edges of $\mathcal{G}_{N}$ are between vertices in consecutive time-steps. For each $k\in\mathbb{Z}_{+}$, there are two edges connecting each of the $K^{(k)}_{N}$ children with its parent(s). If the child has two distinct parent (outcrossing), then there is one edge to each parent; if the child has only one parent (selfing), then there are two edges connecting the child to its parent (hence $\mathcal{G}_{N}$ is a multigraph). For each of the remaining $N-K^{(k)}_{N}$ individuals, there is a single edge connecting to itself in the consecutive time-steps.
The importance of the pedigree, as described in greater detail in the section “Previous Work on Pedigrees” of DFBW (24), is that it encapsulates population dynamics that are common to every single locus on the genome. Even loci extremely far apart are coupled by its dynamics. In particular, gene genealogies far apart on the genome are described by conditionally independent realizations of the ancestral process with respect to the pedigree. Gene genealogies, even conditional on the pedigree, are stochastic due to Mendelian randomness. This is demonstrated in Figure 3. Classical coalescent theory implicitly averages over realizations of the pedigree to determine the average gene genealogy. However, this annealing forces distinct loci to be independent, which is a priori unjustifiable due to the stochasticity of the gene genealogies conditional on the pedigree. Indeed, early biological works of BNA (90); WKLR (12) demonstrate how the pedigree can affect the structure of gene genealogies, and the mathematically rigorous works of Tyu (15); DFBW (24); NWF (25); ABFW (25) have begun to show new scaling limits, and the limitations of the annealing of classical coalescent theory.
Figure 3: (Same pedigree but different genealogies). In the figure we see two different genealogical histories. We focus on a sample of $n=3$ lineages and trace their history backwards in time. Both histories are subject to the same pedigree, that displayed on the right of Figure 2, and yet the history of the sample lineages backwards in time are distinct.
3 Ancestral lines and the coalescent process for a sample
Next, we let $n\in\{2,3,\ldots\}$ be the sample size and consider the lineages of $n$ sampled gene copies under our model, when the population size $N$ is larger than $n$. Suppose we sample $n$ distinct gene copies $\{X_{i}^{N}(0)\}_{1\leq i\leq n}\subset\{0,1\}\times[N]$ from $n$ distinct individuals $\{\hat{X}_{i}^{N}(0)\}_{1\leq i\leq n}\subset[N]$, one gene copy from each individual, at time-step 0 (the present). This assumption can be relaxed, see Remark 4.9.
3.1 The ancestral partition process for a sample of size $n$
For $k\in\mathbb{Z}_{+}$ we let $X_{i}^{N}(k)$ be the gene ancestral to $X_{i}(0)$ $k$ time-steps in the past. Under our model, we can write
$X_{i}^{N}(k):=\left(M_{i}^{N}(k),\hat{X}_{i}^{N}(k)\right)\in\{0,1\}\times[N],\quad i\in[n],$ (3.1)
where $\hat{X}_{i}^{N}(k)$ is the individual in which the gene $X_{i}^{N}(k)$ resides, and
[…]
Given the pedigree between time-steps k1k-1 and $k$, $\hat{X}_{i}^{N}(k)$ is chosen uniformly among the vertices in time-step $k$ to which $\hat{X}_{i}^{N}(k-1)$ is adjacent.
For any given sample of genes at $t=0$ and for any given pedigree, there generally are many possible ways in which to trace these samples backwards in time. This is illustrated in Figure 3, where the (labeled) tree structure of $n=3$ ancestral lines (genetic lineages) backwards in time is evidently not fully determined by the pedigree.
Definition 3.1 (Ancestral line).
For each $1\leq i\leq n$, the process $X_{i}^{N}=(X_{i}^{N}(k))_{k\in\mathbb{Z}_{+}}$ is called the ancestral line of the $i$-th sampled gene $X_{i}(0)$.
Each process $X_{i}^{N}$ is a discrete-time Markov chain taking values in $\{0,1\}\times[N]$. Furthermore, these $n$ processes $\left(X_{i}^{N}\right)_{1\leq i\leq n}$ are correlated Markov chains that form a family of coalescing random walks on $\{0,1\}\times[N]$. In particular, for all $i,j\in[n]$, $X_{i}^{N}(k+1)=X_{j}^{N}(k+1)$ whenever $X_{i}^{N}(k)=X_{j}^{N}(k)$. See Figure 3 for an illustration. If we condition on the pedigree, then these random walks are conditionally independent, and the transition probabilities conditional on the pedigree is given by the Mendelian randomness.
The size of the state space of the family of coalescing random walks, namely $(\{0,1\}\times[N])^{n}$, is of order $N^{n}$ which is huge when $N$ is large. It is customary to reduce the state space by ignoring the indices of the ancestral individuals and only keep track of the coalescence of sample genealogies as we go backward in time. This is done via partitions of $[n]$, or equivalently, equivalence relations on $[n]$. For $1\leqslant i,j\leqslant n$, $k\in\mathbb{Z}_{+}$, we write $i\sim_{k}j$ if and only if samples $i$ and $j$ descend from the same chromosome $k$ time-steps ago, i.e.,
$i\sim_{k}j\quad\Longleftrightarrow\quad X_{i}(k)=X_{j}(k).$ (3.2)
This way, one obtains a stochastic process (called a coalescent process) with state space $\mathcal{E}_{n}$, the space of partitions of $[n]$. The initial state of this coalescent process is the partition into singletons $\xi_{0}^{n}:=\{\{i\}\}_{1\leq i\leq n}$, since we sampled $n$ distinct gene copies, and will end up being the partition into a single block $\mathbf{1}_{n}:=\{[n]\}$ which is the most recent common ancestor (MCRA) of the sample.
For our diploid population, we need to account for which ancestral individuals contain two ancestral gene copies. For this, we use notation from MS (03) and define the ancestral process on the state space
$\mathcal{S}_{n}=\big\{\left\{\left(C_{1},C_{2}\right),\ldots,\left(C_{2x-1},C_{2x}\right),C_{2x+1},\ldots,C_{b}\right\}:b\in[n],1\leqslant x\in\lfloor{b}/{2}\rfloor,\{C_{1},\ldots,C_{b}\}\in\mathcal{E}_{n}\big\},$
where $\lfloor x\rfloor$ is the largest integer less than or equal to $x$. We equip both spaces $\mathcal{E}_{n}$ and $\mathcal{S}_{n}$ with the discrete topology. Hence, each element $\xi\in\mathcal{S}_{n}$ is of the form
$\{(C_{1},C_{2}),(C_{3},C_{4}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\},$ (3.3)
[…]
Definition 3.2.
We define an $\mathcal{S}_{n}$-valued stochastic process $\chi^{N,n}=(\chi^{N,n}(k))_{k\in\mathbb{Z}_{+}}$ as follows: for $k\in\mathbb{Z}_{+}$,
  • $i$ and $j$ are in the sample block in $\chi^{N,n}(k)$ if and only if $X_{i}(k)=X_{j}(k)$, i.e. the ancestral lines of the $i$-th and the $j$-th samples coalesced $k$ time-steps in the past, and
  • two blocks are in a set together in $\chi^{N,n}(k)$ if and only if $\hat{X}_{i}(k)=\hat{X}_{j}(k)$, i.e. the two extant lineages corresponding to these blocks are in the same individual in the population $k$ time-steps in the past.
We call this process $\chi^{N,n}$ the ancestral process of the sample.
We will investigate the convergence of a time-rescaling of $\chi^{N,n}$ with a random time change $S$ defined by
$S(k):=\sup\left\{l\in\{0,1,2,\ldots,k\}:\,\chi^{N,n}(l)\in\mathcal{E}_{n}\right\},$ (3.4)
the last time-step (up to $k$) at which the ancestral process $\chi^{N,n}$ is totally dispersed (i.e. lies in $\mathcal{E}_{n}$). Note that $S(k)$ is well-defined since by our assumption $\chi^{N,n}(0)\in\mathcal{E}_{n}$. We define $\widetilde{\chi}^{N,n}$ to be the random time-change of $\chi^{N,n}$ as follows:
$\widetilde{\chi}^{N,n}(k):=\chi^{N,n}(S(k)).$
The $\mathcal{E}_{n}$-valued process $\widetilde{\chi}^{N,n}$ is non-Markovian. However, the random time-change (3.4) not only keeps the process to live in the state space $\mathcal{E}_{n}$ for which our weak convergence method works, but also prevents accumulation of jumps of the process in the limit as the population size $N$ goes to infinity.
3.2 An annealed scaling limit
To connect to classical results in coalescent theory, including Kin (82); Möh (98); Pit (99); Sag (99); Sch (00); BLS (18), we shall consider the unconditional (a.k.a. annealed) law, one that averages over the random pedigree, of the ancestral process.
Denote by
$\mathfrak{p}_{\xi\eta}^{N,n}:=\mathbb{P}\left(\chi^{N,n}(1)=\eta\mid\chi^{N,n}(0)=\xi\right),$ (3.5)
the one-step transition probabilities for the ancestral process $\chi^{N,n}$. Let $c_{N}=\mathfrak{p}_{\xi_{0}^{2}\mathbf{1}_{2}}^{N,n}$ be the probability that two randomly chosen genes from two distinct individuals coalesce in one time-step in the past. Then
$c_{N}=$ $\mathbb{E}\left[\frac{(N-K)K}{N^{2}\,(N-1)}\right]\;+\;\frac{1}{\binom{N}{2}}\,\mathbb{E}\!\left[\sum_{i=1}^{N}\!\left(\frac{1}{2}\binom{V_{i,i}}{2}+\frac{1}{4}\,V_{i,i}\!\sum_{j\neq i}V_{i,j}+\frac{1}{8}\binom{\sum_{j\neq i}V_{i,j}}{2}\right)\right$]
$=$ $\mathbb{E}\left[\frac{(N-K)K}{N^{2}\,(N-1)}\right]\;+\;\frac{1}{\binom{N}{2}}\,\mathbb{E}\!\left[\frac{1}{16}\sum_{i=1}^{N}\bigl(\tilde{V}_{i}^{\,2}-\tilde{V}_{i}\bigr)-\frac{S}{8}\right],$ (3.6)
where the first term comes from one newborn and one carryover, and the second term comes from two newborns. This formula generalizes that in (BLS, 18, Equation 1.4).
Convergence will be established for the time-rescaled ancestral process $\left(\bar{\chi}^{N,n}(t)\right)_{t\in\mathbb{R}_{+}}$ defined by
$\bar{\chi}^{N,n}(t)=\widetilde{\chi}^{N,n}\left(\lfloor tc_{N}^{-1}\rfloor\right)=\chi^{N,n}\left(S\lfloor tc_{N}^{-1}\rfloor\right).$ (3.7)
We can and will consider this process $\bar{\chi}^{N,n}=\left(\bar{\chi}^{N,n}(t)\right)_{t\in\mathbb{R}_{+}}$ as a random variable taking value in the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$ equipped with the $J_{1}$ topology (see EK (09)).
We introduce three assumptions on the asymptotic behaviors of our model as $N\to\infty$ below. The first assumption ensures a continuous-time limiting model.
Assumption 3.3.
Suppose $\lim_{N\to\infty}c_{N}=0$, or equivalently, $\lim_{N\to\infty}\frac{\mathbb{E}[V_{1}^{2}]}{N}=0$ where $V_{i}:=\sum_{j=1}^{N}V_{i,j}$ is the total number of offspring for individual $i$.
Asymptotically, one unit of (continuous) time corresponds to $\lfloor c_{N}^{-1}\rfloor$ time-steps. To quantify the time-scale on which distinct sample lineages in the same individual disperse, we let $d_{N}$ denote the probability that, in a single time-step in the past, two sample lineages in the same individual would disperse into two different individuals. Precisely,
$d_{N}:=\mathbb{P}\left(\chi^{N,2}(1)=\Big\{\{1\},\{2\}\Big\}\mid\chi^{N,2}(0)=\Big\{(\{1\},\{2\})\Big\}\right)=\mathbb{E}\left[\frac{K_{N}-S_{N}}{N}\right]=\frac{1-\alpha_{N}}{N}\mathbb{E}\left[K_{N}\right].$ (3.8)
We refer to $d_{N}^{-1}$ as the dispersal or outcrossing timescale.
The present work focuses on the asymptotic regime where the outcrossing timescale is comparable to, or slower than, the coalescence timescale. We refer this as the limited outcrossing regime, which generalizes the setting of the Moran model studied in NWF (25). This regime is formalized in Assumption 3.4 below.
Assumption 3.4 (Limited outcrossing).
We assume that $\alpha_{N}\to 1$ and $d_{N}c_{N}^{-1}\to\lambda\in\mathbb{R}_{+}$ as $N\to\infty$.
Assumption 3.4 together with (3.8) implies that
$1-\alpha_{N}\;=\;\frac{N\,c_{N}}{\mathbb{E}[K_{N}]}\,\lambda\,(1+o(1))\to 0$ (3.9)
as $N\to\infty$. Hence, $\alpha_{N}\to 1$ at a rate proportional to $\frac{N\,c_{N}}{\mathbb{E}[K_{N}]}$ when $\lambda>0$,
Remark 3.5.
It is notable that the assumption $d_{N}c_{N}^{-1}\to\lambda\in\mathbb{R}_{+}$ together with Assumption 3.3 implies that the selfing probability $\alpha_{N}$ tends to 1 in many cases including the case when $K_{N}=N$ (e.g. Wright-Fisher model) and $K_{N}=O(1)$ (e.g. Moran model). Indeed, (3.8) and the assumption $d_{N}c_{N}^{-1}\to\lambda$ imply the equality in (3.9). Assumption 3.3 then implies that $\alpha_{N}\to 1$ if $\mathbb{E}[K]$ is of order $N$. On other hand, note that (3.6) and the fact $\sum_{i=1}^{N}\tilde{V}_{i}=2K$ implies that $c_{N}\leq C\,\frac{\mathbb{E}[K+K^{2}]}{N^{2}}$ for some constant $C$ independent of $N$. Hence $\alpha_{N}\to 1$ if $\frac{\mathbb{E}[K^{2}]}{N\mathbb{E}[K]}\to 0$.
Our third and last assumption is on the generator of the ancestral process. Let $F_{\rm hap}:\mathcal{S}_{n}\to\mathcal{E}_{n}$ denote the haploid map defined by
$\{(C_{1},C_{2}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\}\mapsto\{C_{1}\cup C_{2},\ldots,C_{2x-1}\cup C_{2x},C_{2x+1},\ldots,C_{b}\}.$ (3.10)
This map announces that lineages in the same individuals have coalesced, essentially reducing a diploid model to a haploid model. This map is different from the “complete dispersal” map ${\rm cd}:\mathcal{S}_{n}\to\mathcal{E}_{n}$ used in BLS (18); ABFW (25).
Define the matrix $\mathit{P}_{N,n}=\left(\mathit{p}_{\xi\eta}^{N}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ by
$\mathit{p}_{\xi\eta}^{N}:=\mathbb{P}\left(F_{\rm triv}\left(\chi^{N,n}(1)\right)=\eta\,\Big|\,\chi^{N,n}(0)=\xi\right)=\sum_{\zeta\in\mathcal{S}_{n}:\,F_{\rm triv}(\zeta)=\eta}\mathfrak{p}_{\xi\zeta}^{N}.$
Assumption 3.6.
For any integer $n\geq 2$, the limit
$Q_{n}:=\lim_{N\to\infty}c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)$
exists as a real-valued matrix $Q_{n}=(q^{n}(\xi,\eta))_{\xi,\eta\in\mathcal{E}_{n}}$. Furthermore, $\left(Q_{n}\right)_{n\geq 2}$ is a consistent family, which means that for any $m<n$ the pushforward of $Q_{n}$ under $\varpi_{n,m}$ is $Q_{m}$ (i.e. $q^{n}(\xi,\eta)=q^{m}\big(\varpi_{n,m}(\xi),\varpi_{n,m}(\eta)\big)$ for all $\xi,\eta\in\mathcal{E}_{n}$), where $\varpi_{n,m}:\mathcal{E}_{n}\to\mathcal{E}_{m}$ is the projection obtained by restricting the partition to $[m]$.
To describe the annealed limit, we also describe an $n$-$\Xi$-coalescent following (BLS, 18, P.4).
Definition 3.7.
Let $\Xi$ be a finite measure on the infinite ordered simplex
$\Delta:=\{x=(x_{1},x_{2},\ldots)\in[0,1]^{\infty}:x_{1}\geq x_{2}\geq\ldots,\sum_{i=1}^{\infty}x_{i}\leq 1\}.$
An $n$-$\Xi$-coalescent is a continuous-time Markov process $\chi^{n}=\left(\chi^{n}(t)\right)_{t\in\mathbb{R}_{+}}$ taking values in $\mathcal{E}_{n}$, whose transition rates are invariant under permutation and are described as follows: For any $\xi$ in $\mathcal{E}_{n}$ of consisting of $b$ blocks, suppose that $\eta$ is an element of $\mathcal{E}_{n}$ obtained from $\xi$ by keeping $s$ of the blocks of $\xi$ the same, and coalescing the remaining bsb-s blocks into $r$ blocks of sizes $k_{1},k_{2},\ldots,k_{r}.$ The transition rate from $\xi$ to $\eta$ is given by
$\lambda_{b;k_{1},\ldots,k_{r};s}:=\mathbbm{1}_{(r,s)=(1,b-2)}\Xi(\mathbf{0})+\int_{\Delta\setminus\{\mathbf{0}\}}\sum_{l=0}^{s}\sum_{i_{1},\ldots,i_{r+l}\text{ distinct}}^{\infty}\binom{s}{l}\prod_{m=1}^{r}x_{i_{m}}^{k_{m}}\prod_{m=r+1}^{r+l}x_{i_{m}}(1-\left|x\right|)^{s-l}\frac{\Xi(dx)}{\left\langle x,x\right\rangle},$
where $\left\langle x,x\right\rangle=\sum_{i}x_{i}^{2}$ and $\left|x\right|=\sum_{i}x_{i}$.
Informally, an $n$-$\Xi$-coalescent can be described in terms of interval partitions. Take $x\in\Delta\setminus\{\mathbf{0}\}$ at rate $\frac{1}{\left\langle x,x\right\rangle}\Xi$. As $\sum_{i}x_{i}\leq 1$, $x$ decides an interval partition. We throw each block of the coalescent uniformly at random on the unit interval. If they belong to the same element of the partition, and if they are at most $\sum_{i}x_{i}$, then they coalesce. Additionally, each pair of blocks coalesces independently at rate $\Xi(\mathbf{0})$.
Remark 3.8.
We shall show in Lemma Appendix A.1 that, under Assumption 3.3 and a usual condition on the ordered offspring distribution, we have that Assumption 3.6 holds and that for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent; i.e. $\lambda_{b;k_{1},\ldots,k_{r};s}=q^{n}(\xi,\eta)$ whenever $\xi,\eta\in\mathcal{E}_{n}$ and $\lambda_{b;k_{1},\ldots,k_{r};s}$ are as in Definition 3.7.
Assumption 3.6 implies that there is a unique finite measure $\Xi$ on $\Delta$ for which $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent governed by $\Xi$ for all $n\geq 2$. This is precisely the content of (Ber, 04, Proposition 3), which shows that the generator $C$ of the coalescence part of an exchangeable fragmentation-coalescence (EFC) process is completely determined by the consistent family $\left(Q_{n}\right)$ of its finite-dimensional restrictions, each $Q_{n}$ encoding the jump rates on $\mathcal{E}_{n}$. Moreover, (Ber, 04, Proposition 3) proves that any such consistent and exchangeable collection of generators $\left(Q_{n}\right)_{n\geq 2}$ arises as the finite-dimensional projections of an EFC process taking values in $\mathcal{E}_{\infty}$. In the purely coalescent case, this implies that any consistent family of $n$-$\Xi$-coalescents can be realized as the projections of an infinite $\mathcal{E}_{\infty}$-valued Markov process with infinitesimal generator $Q$ determined by $\Xi$.
We can now describe the convergence of the unconditional law of the time-rescaled ancestral process $\left(\bar{\chi}^{N,n}(t)\right)_{t\in\mathbb{R}_{+}}$ defined by (3.7), under the model presented in Section 2.
Theorem 3.9 (Annealed convergence).
Suppose that Assumptions 3.3, 3.4 and 3.6 hold. Then there is a unique finite measure $\Xi$ on $\Delta$, made precise in Remark 3.8, such that $\bar{\chi}^{N,n}$, as a $\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$-valued random variable, converges in law to an $n$-$\Xi$-coalescent governed by $\Xi$ as $N\to\infty$.
Remark 3.10.
This result complements that given in BLS (18) by providing the high-selfing analogue of Theorem 1.1 contained therein. That Theorem 3.9 is not extended to the case $d_{N}c_{N}^{-1}\to\infty$ is not due to any particular difficulty with the annealed proof. Rather, the tools for the quenched proof for the the case $d_{N}c_{N}^{-1}\to\infty$ are sufficiently different from those required for the present work.
The proof of Theorem 3.9 is postponed to Section Appendix B, as it will follow from results in Section 4.
4 Quenched convergence of the ancestral process
Recall the pedigree $\mathcal{G}_{N}$ from Definition 2.3 and define the $\sigma$-algebra $\mathcal{A}_{N}$ generated by the pedigree $\mathcal{G}_{N}$ and the labels of the $n$ distinct individuals from whom we have sampled, i.e.
$\mathcal{A}_{N}:=\sigma(\mathcal{G}_{N},\hat{X}_{i}^{N}(0):1\leq i\leq n).$
We denote the quenched law of the time-rescaled ancestral process (3.7) by
$\mathcal{L}^{N,n}:=\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\mid\mathcal{A}_{N}\right).$ (4.1)
Our main result in this paper, Theorem 4.7, asserts that this $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$-valued random variable converges as $N\to\infty$. Furthermore, the limit is the random law of a family of coalescing random walk on a random graph described by an exchangeable fragmentation-coalescence processes (EFCs). The latter was introduced in Ber (04) which we describe in the next section, before giving the rigorous statement of our main result in Theorem 4.7.
Briefly, let $\lambda$ and $Q$ be given by Assumption 3.4 and Remark 3.8 respectively. Then $\mathcal{L}^{N,n}$ converges in distribution to a random element $\mathcal{L}_{Q,\lambda}^{n}\in\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ described as follows: Consider a random graph $G_{Q,\lambda}^{n}=\big(G_{Q,\lambda}^{n}(t)\big)_{t\in\mathbb{R}_{+}}$, to be called a $Q$-$\lambda$ graph starting with $n$ nodes at $t=0$, in which each node in the graph splits into two nodes independently at rate $\lambda$, and where any subset of nodes coagulate into one node according to the rate matrix $Q$ (to be made precise in Definition 7.1); see Figure 4 for an illustration. Given $G_{Q,\lambda}^{n}$, we consider a system of coalescing random walks on it, where each random walker chooses one of the two directions to go when it reaches a point of fragmentation. Then $\mathcal{L}_{Q,\lambda}^{n}$ is the conditional law of this system of coalescing random walks given $G_{Q,\lambda}^{n}$; we make this precise in Definition 7.3.
4.1 Random walks on exchangeable fragmentation-coalescence processes
The (random) $Q$-$\lambda$ graph $G_{Q,\lambda}^{n}$ mentioned above is a subgraph of an exchangeable fragmentation-coalescence process (EFC), where the EFC was introduced in Ber (04). Here we give a concise description of EFCs.
Definition 4.1 (Ber (04)).
An exchangeable fragmentation-coalescence (EFC) process $\Pi$ is a $\mathcal{E}_{\infty}$-valued exchangeable Markov process so that the restriction to $\mathcal{E}_{n}$, denoted by $\Pi^{n}$, is a cádlág finite state Markov chain which can only evolve by fragmentation of one block or by coalescence.
By Ber (04)[p. 781] $\Pi$ is fully characterized by $\Pi(0)$, finite measures $\nu_{Disl}$ and $\nu_{Coag}$ on $\Delta$, and non-negative constants $c_{k},c_{e}$. In this paper we will always consider the case $\nu_{Disl}=0$. $c_{k}$ and $\nu_{Coag}$ form the coalescing part of the EFC. Specifically, for any finite subset $l$ of particles in the EFC, they coalesce as if they were an $l$-$\Xi$-coalescent where $\Xi=c_{k}\delta_{\mathbf{0}}+\nu_{Coag}$. That is, the non-Kingman coalescences are governed by $\nu_{Coag}$ and $c_{k}$ is the rate at which any pair of particles coalesce. $c_{e}$ is simply the rate at which each particle in the EFC fragments into two particles.
Example 4.2 (Ancestral Recombination Graph).
Suppose $\Pi=\left(\Pi(t)\right)_{t\in\mathbb{R}_{+}}$ is an EFC with characteristics $c_{k}=2$, $\nu_{Coag}=0$, $c_{e}=\lambda,$ and $\nu_{Disl}=0$. Then for any $n$ particles, each of the $\binom{n}{2}$ pairs of particles coalesce independently with rate $2$ and each of the $n$ particles fragment independently at rate $\lambda$, which is preceisely the structure of an ancestral selection or ancestral recombination graph Gri (91); GM (97); KN (97). This subgraph of $\Pi$ was proposed in (NWF, 25, Section 4.1) as the scaling limit of a diploid Moran model with selfing. This is established in Section 5.1.
Example 4.3.
Suppose that $\psi$ is a fixed constant in $[0,1]$. By a $(\psi,\lambda,\rho)$-EFC $\Pi(\psi,\lambda,\rho)$ we mean an EFC where $c_{k}=2$, $c_{e}=\lambda$, and
$\nu_{Coag}=\frac{\psi^{2}}{2}\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,0,\ldots\right)}.$
This EFC has implicitly appeared in BBE (13) when $\psi$ is a fixed constant.
The example 4.3 generalizes when we allow $\psi$ to be random quantity. A natural prior is to take $\psi$ to be Beta-distributed. This yields an EFC whose coalescent part is a beta coalescent.
Example 4.4.
Suppose that $\psi$ is random and Beta(2r,r)(2-r,r) distributed for some $r\in[0,1]$. Let $\nu_{Coag}$ denote the measure on $\Delta\setminus\{\mathbf{0}\}$ defined by
$\int_{0}^{1}\frac{z^{2}}{2}\delta_{\left(\frac{z}{2},\frac{z}{2},0,0,\ldots\right)}\,{\rm Beta}(2-r,r)(dz),,$
where the ${\rm Beta}(2-r,r)(dz)$ measure has density
$\frac{1}{\Gamma(2-r)\Gamma(r)}z^{1-r}(1-z)^{r-1},\quad z\in(0,1).$
There is a three-parameter family of EFCs $\Pi(r,\rho,\lambda)$ with characteristics $c_{k}=2$, $\nu_{\rm Coag}=\rho\Xi^{\prime}(\psi)$, $c_{e}=\lambda$, and $\nu_{\rm Frag}=0$ that arise as the scaling limit of models where a highly reproductive couple gives birth to $\psi\simBeta$(2r,r)(2-r,r) of the population, as in BBE (13). This is established in Section 5.4.
Remark 4.5.
The ancestral recombination graph of BBE (13) is in fact the subgraph of the $\Pi$ of Example 4.3 given by tracing all trajectories of the first $n$ blocks of $\Pi$ backwards in time.
We define now a family of coalescing random walks $\left(x_{i}\right)_{i=1}^{n}$ on an EFC $\Pi$ with $\Pi(0)=\xi_{0}$. Two realizations of these coalescing random walks on a given $\Pi$ are demonstrated in Figure 4. We follow the Poissonian construction of Ber (04). Let $P_{C}=\left(\left(t,\xi^{(C)}(t)\right)\right)_{t\geq 0}$ and $P_{F}=\left((t,\xi^{(F)}(t),k(t))\right)_{t\geq 0}$ be two independent Poisson point processes (PPPs) on the same filtration. The atoms of $P_{C}$ are points in $\mathbb{R}_{+}\times\mathcal{E}_{\infty}$ and $P_{C}$ has intensity measure $dt\otimes C$. The atoms of $P_{F}$ are points in $\mathbb{R}_{+}\times\mathcal{E}_{\infty}\times\mathbb{N}$ and $P_{F}$ has intensity measure $dt\otimes F\otimes\#$, where #\# is the counting measure on $\mathbb{N}$. $\Pi$ may be constructed in the obvious way from $P_{C}$ and $P_{F}$, as described in (Ber, 04, Section 3.2). We addend to this description a construction of coalescing random walks $\left(x_{i}\right)$ on top of $\Pi$. The $\left(x_{i}\right)$ are a family of $n$ coalescing particles $x_{i}$, modeled as $\mathbb{N}$-valued random variables, tracking the label of the block in $\Pi$ to which the $i$th particle belongs. We restrict ourselves to the case where the fragmentation $F=\lambda\mathbf{e}$, which is the case relevant to our results. In this case, the fragmentation of a block fragments it into precisely two parts.
Figure 4: Here we see on a fixed realization of an EFC $\Pi$ with $c_{k}=1$ two realizations of coalescing random walks on $\Pi$. These random walks follow each coalescence in the EFC and choose between each of the possible edges ahead of them at a fragmentation with equal likelihood.
Let $m(t):=\max_{i}x_{i}(t)$ We begin with $x_{i}(0)=i$ for each $1\leq i\leq n$ and construct $x_{i}$ as follows:
  • if $t$ is not an atom time for either $P_{C}$ or $P_{F}$ then $x_{i}(t-)=x_{i}(t)$,
  • if $t$ is an atom time for $P_{C}$ such that $\xi^{(C)}(t)_{\mid m(t-)}\neq\xi_{0}^{m(t-)}$ then $x_{i}(t)$ is equal to the label of the block to which the $x_{i}(t-)$th block of $\Pi^{m(t-)}$ is mapped,
  • if $t$ is an atom time for $P_{F}$ such that $k(t)=x_{i}(t-)$ then the label in $\Pi(t)$ of one of the two blocks into which the $x_{i}(t-)$th block fragments is chosen fairly (i.e. probability $\frac{1}{2}$ each) and $x_{i}(t)$ is set equal to this label for all $i$ with $x_{i}(t-)=k(t)$.
Definition 4.6.
The coalescent process $\chi^{n}$ defined by
$i\sim_{\chi^{n}(t)}j\text{ if }x_{i}(t)=x_{j}(t)$
is a $\mathcal{E}_{n}$-valued process. We denote by $\mathcal{L}_{\Pi}^{n}\in\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ the conditional law of $\chi^{n}$ given $\Pi$.
We provide, in the main result of this paper contained in Section 4.2 criteria by which $\mathcal{L}_{\Pi}^{n}$ is the quenched scaling limit of the law of the coalescent conditional on the pedigree.
4.2 Main result
We are now ready to state our main result in this paper. Recall that $\mathcal{L}^{N,n}:=\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\mid\mathcal{A}_{N}\right)$ was defined in (4.1), and $\mathcal{L}_{\Pi}^{n}$ was defined in Definition 4.6.
Theorem 4.7 (Quenched convergence).
Suppose Assumptions 3.3, 3.6 and 3.4 hold. Then there is a unique measure $\Xi$ on $\Delta$, as explained in Remark 3.8 such that, for $\Pi$ an EFC with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi-c_{k}\delta_{\mathbf{0}}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$, it holds that $\mathcal{L}^{N,n}$ converges in distribution to $\mathcal{L}_{\Pi}^{n}$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ as $N\to\infty$.
Theorem 4.7 says that when the outcrossing timescale and the coalescence timescale are comparable, the structure of the $n$ sample lineages, conditional on the pedigree, converges to that of coalescing random walks on a realization of an EFC process.
Remark 4.8.
Note that the form of weak convergence in distribution is stronger than that described by the Meyer-Zheng topology of MZ (84). In particular, it follows from Theorem 4.7 that the random occupation measure in $\mathcal{M}_{\rm loc}\left(\mathbb{R}_{+}\times\mathcal{E}_{n}\right)$ defined by $dt\otimes\delta_{\bar{\chi}^{N,n}(t)}$ converges vaguely in distribution with respect to the measure $\mathbb{P}\left(\cdot\,\mid\,\mathcal{A}_{N}\right)$ to the random measure $dt\otimes\delta_{\chi^{n}(t)}$ governed by $\mathbb{P}\left(\cdot\,\mid\,\Pi\right)$.
Remark 4.9.
[Relaxing assumption on sampling]In our main result, we assumed that the $n$ sampled gene copies are from $n$ distinct individuals, one gene copy from each individual. This assumption can be relaxed and the corresponding result can be obtained by our method, and we briefly describe the corresponding limiting object here. Suppose the $n$ sampled gene copies consist of $2m$ gene copies sampled in pairs within individuals and n2mn-2m gene copies sampled each from distinct individuals. Then, since $\alpha_{N}\to 1$, the two gene copies within each of the $m$ pairs will instantaneously coalesce with probability 1. Both the annealed limit and the quenched limit (as $N\to\infty$) will be the same as before, but starts with nmn-m gene copies rather than $n$. The annealed limit in this case (for fixed selfing probability) was obtained in ND (97). For the proof, one may specify what it means to take supremum in an empty set in 3.4 and show that the transient behavior of the discrete-time processes in $\mathcal{S}_{n}$ converge to that of the limiting process.
5 Applications
In this section, we discuss some important applications of our main result that motivate this study.
In Section 5.1 we establish that the coalescing random walk on the classical ancestral recombination graph (ARG), first introduced for a diploid Moran model with selfing in (NWF, 25, Section 4.1), arises as the limiting conditional coalescent for the corresponding diploid Wright-Fisher model, and to perturbations of both of these models as well.
In Section 5.2, we prove that the form of convergence for $\mathcal{L}^{N,n}$ implies the convergence of important tree statistics in population genetics, in particular the branch lengths associated with the site-frequency spectrum, or SFS. We also illustrate how the SFS depends on the pedigree when the selfing probability is close to one.
In Section 5.2, we establish that some important tree statistics in population genetics converge. Of particular importance in population genetics are certain classes of integral functionals, namely the internal and external branch lengths of the coalescent process. These statistics form the basis of inference methods in population genetics, and so we need to know that the form of convergence for $\mathcal{L}^{N,n}$ implies the convergence of these statistics. The adequate notion of convergence necessary is weak convergence in distribution of random measures, which we make precise in Section 6. The standard reference is (Kal, 17, Chapter 4).
We then provide applications of Theorem 4.7 to our diploid Sargasyan-Wakeley model in Section 5.3, and examples of how one can establish scaling limits for models like those in BBE (13); DFBW (24) given by mixed demographies in Section 5.4.
5.1 Robustness of coalescing walks on the ancestral recombination graph
In this section we establish, as a corollary of Theorem 4.7 that the scaling limit described in (NWF, 25, Section 4.1) is robust to perturbations of the pedigree structure. That is, so long as non-binary mergers are vanishingly unlikely on the coalescent time-scale, the scaling limit is given by coalescing random walks on EFCs of the form described in Example 4.2.
Let $\mathfrak{c}_{3}:=\mathbb{P}_{\xi_{0}^{3}}\left(\chi^{N,n}(1)=\mathbf{1}_{3}\right)$ be the probability that three distinct sample lineages in three distinct individuals coalesce in a single time-step. This is the simplest type of non-binary merger. Then
$\mathfrak{c}_{3}=$ $\frac{1}{\binom{N}{3}}\,\mathbb{E}\bigg[\frac{N-K}{N}\sum_{i=1}^{N}\!\left(\frac{1}{4}\binom{V_{i,i}}{2}+\frac{1}{8}\,V_{i,i}\sum_{j\neq i}V_{i,j}+\frac{1}{16}\binom{\sum_{j\neq i}V_{i,j}}{2}\right)$
$+$ $\sum_{i=1}^{N}\!\left(\frac{1}{8}\binom{V_{i,i}}{3}+\frac{1}{16}\binom{V_{i,i}}{2}\,\sum_{j\neq i}V_{i,j}+\frac{1}{32}\,V_{i,i}\binom{\sum_{j\neq i}V_{i,j}}{2}+\frac{1}{64}\binom{\sum_{j\neq i}V_{i,j}}{3}\right)\bigg$], (5.1)
where, in the first equality, the first term comes from two newborns and one carried-over (parent $i$, the common parent for all 3 lineages) in which both newborn transmissions must pick the carried-over lineage’s copy in $i$; and the second term comes from three newborns in which all three transmissions through the same parent $i$ must choose the same copy. The proof of Equation (5.1) is given in Lemma Appendix C.1 in the Appendix.
Assumption 5.1 (Negligible triple coalescent).
We assume that $\mathfrak{c}_{3}=o(c_{N})$, i.e. $\mathfrak{c}_{3}c_{N}^{-1}\to 0$ as $N\to\infty$.
Theorem 5.2.
Suppose that Assumptions 3.3, 3.4 and 5.1 hold. Then Assumption 3.6 holds and the sequence of random measures $\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ converge weakly in distribution to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is the ancestral recombination graph described in Example 4.2.
Proof.
By Lemma Appendix C.2, $p_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$. Therefore the limiting exit rate from the state $\xi$ is $2\binom{\xi}{2}$ when we take the $c_{N}$ time-rescaling with $c_{N}\to 0$. Further, Lemma Appendix C.2 demonstrates that $\mathit{p}_{\xi\eta}^{N,n}\in o(c_{N})$ for any $\eta$ that cannot be obtained from $\xi$ via a single binary merger, i.e. for which $\xi\not\prec\eta$. For such $\xi,\eta$ we have that the limiting exit rate from $\xi$ to $\eta$ is 0 with the $c_{N}$ time-rescaling with $c_{N}\to 0$. Finally, there are $\binom{|\xi|}{2}$ partitions $\eta$ for which $\xi\prec\eta$. By exchangeability, these are equally likely and so the limiting transition rate from $\xi$ to $\eta$ is $2$ when $\xi\prec\eta$.
It follows from the above that $Q_{n}$ from Assumption 3.6 exists and is given by
$Q_{n}(\xi,\eta)=\begin{cases}-2\binom{|\xi|}{2}&,\text{ if }\xi=\eta\\ 2&,\text{ if }\xi\prec\eta\\ 0&,\text{ otherwise}\end{cases}.$
$Q_{n}$ is the transition-rate matrix of a Kingman $n$-coalescent with time rescaled by a factor of $2$. Hence Assumption 3.6 is satisfied and $(Q_{n})_{n\in\mathbb{N}}$ determines a measure $\Xi=2\delta_{\mathbf{0}}$ on $\Delta$ in the sense of Remark 3.8. As Assumptions 3.3 and 3.4 hold, and we have shown that $Q_{N,n}$ converges to the generator of a time-rescaled Kingman $n$-coalescent, the claim follows directly from Theorem 4.7. ∎
Remark 5.3.
Theorem 5.2 establishes scaling limits for the models of ND (97); Möh (98); NWF (25), which are alike insofar as they all have a time-rescaled Kingman coalescent for their annealed limits. The models of Möh (98) and ND (97) are diploid Wright-Fisher models, i.e. $K_{N}=P_{N}=N$, while the model of NWF (25) is diploid Moran model, i.e. $K_{N}=1$ and $P_{N}=N$. This corollary establishes that the model of (NWF, 25, Section 4.1) is robust as long as the annealed limit of the models is Kingman-like with high selfing rate.
5.2 Convergence of tree statistics
The notion of convergence in what follows is weak convergence in distribution of random measures, which we make precise in Section 6. The standard reference is (Kal, 17, Chapter 4). As described briefly in Remark 4.8, the form of weak convergence established in Theorem 4.7 implies that, for any compactly supported function $\varphi:\mathbb{R}_{+}\times\mathcal{E}_{n}\to\mathbb{R}$, the associated integral functional $\int_{0}^{\infty}\varphi(s,\cdot(s))ds$ converges in distribution, i.e.
$\mathbb{E}\left[\int_{0}^{\infty}\varphi\left(s,\bar{\chi}^{N,n}(s)\right)ds\in\cdot\,\mid\,\mathcal{A}_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\int_{0}^{\infty}\varphi\left(s,\chi^{n}(s)\right)ds\in\cdot\,\mid\,\Pi\right].$
An important class of these integral functionals, which in fact motivated our choice of the form of convergence of the random measures $\mathcal{L}^{N,n}$, are the internal and external branch lengths of the coalescent associated with the SFS. Formally, the $r$th total branch length $\tau^{n,r}$ of a coalescent $\chi^{n}\in\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$ of a sample of size $n$ is defined by
$\tau^{n,r}(\chi^{n})=\int_{0}^{\infty}\#\{C\in\chi^{n}(s):|C|=r\}ds.$
That is, $\tau^{n,r}$ is the sum of the lengths of the edges in the coalescent ancestral to $r$ samples. Note that $\tau^{n,r}$ is not of the form of $\varphi$ described above, so it requires an additional argument to show that these functionals do indeed converge. This is done via a truncation argument.
Proposition 5.4.
Suppose that, as $N\to\infty$, that Assumptions 3.3, 3.4, and 3.6 hold. Then, for any finite $s\in\mathbb{N}$,
$\mathbb{E}\left[\left(\tau^{n,r}(\bar{\chi}^{N,n})\right)_{1\leq r\leq s}\,\mid\,\mathcal{A}_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\tau^{n,r}(\bar{\chi}^{N,n})_{1\leq r\leq s}\,\mid\,\Pi\right],$
where $\Pi$ is as described in Theorem 4.7.
Proof.
For each $r\geq 1$ and $\xi\in\mathcal{E}_{n}$, set $g_{r}(\chi^{n})=\#\{C\in\chi:|C|=r\}$, which is bounded above by $\lfloor n/r\rfloor$. For $T>0$, define the truncated branch length
$\tau^{n,r,T}(\chi)=\int_{0}^{T}g_{r}(\chi(t))\,dt.$
By Remark 4.8, the occupation measures $\mu_{\bar{\chi}^{N,n}}$ converge vaguely to $\mu_{\chi^{n}}$, so for any bounded continuous cutoff $\psi_{T,\varepsilon}\in C_{c}(\mathbb{R}_{+})$ with $\int_{0}^{\infty}|\psi_{T,\varepsilon}(s)-\mathbf{1}_{[0,T]}(s)|ds\leq s$,
$\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\bar{\chi}^{N,n}(t))\,dt\,\mid\,\mathcal{A}_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\chi^{n}(t))\,dt\,\mid\,Pi\right].$
By letting $\varepsilon\to 0$ we obtain $\tau^{n,r,T}(\bar{\chi}^{N,n})\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\tau^{n,r,T}(\chi^{n}).$ Finally, monotone convergence as $T\to\infty$ yields joint convergence of $(\tau^{n,r})_{1\leq r\leq s}$ under the quenched laws. ∎
5.2.1 The site-frequency spectrum under nearly complete selfing
In light of the importance of the SFS and other measures of genetic variation for inferring past events and processes affecting populations, it would be of some interest to go beyond the convergence result in Proposition 5.4 and to describe expected patterns of variation for a range of specific diploid population models conditional on the pedigree. We leave this to future research, as it is outside the scope of the present work, but we note that a key question is the extent to which the SFS and other measures of genetic variation depend on the pedigree.
Here we illustrate this using simulations of the (robust) conditional limit described in Theorem 5.2, namely, the coalescing random walk on the ancestral recombination graph (ARG). In this limit, the ARG representing all the information of the pedigree relevant to the sample is generated by an EFC in which every distinct pair of ancestral lineages coalesce with rate $2$ and every single lineage split at rate $\lambda$. After an exponentially distributed waiting time with rate parameter $n(\lambda+n-1)$, one of these $\binom{n}{2}+n$ events is chosen in proportion to its rate. If two lineages coalesce, the number of ancestral lineages increases by one. If a lineage splits, the number of ancestral lineages increases by one. All events and their times are recorded, and the process stops the first time there is just a single ancestral lineage.
Note that the marginal coalescent process for the sample is a Kingman coalescent process, which does not depend on $\lambda$. However, the time it takes to reach a single lineage in the EFC process when $\lambda$ is large may far exceed any reasonable time for any coalescing random walk on the resulting ancestral graph to reach its most recent common ancestor. Thus, there should be little error in stopping the EFC process before it reaches a single lineage. The way time is measured here, there would be only a $e^{-20}\approx 2\times 10^{-9}$ chance of a pairwise coalescence time greater than $10$ in the marginal process. In our simulations, we set the splitting rate to zero at $10$ units of time, so the remainder of the graph is a single Kingman coalescent tree, which we note preserves the marginal distribution of the process.
The gene genealogy at a locus is generated by tracing ancestral lineages backward in time through the graph, as in the presentation of Figure 4. When a lineage encounters a split, it follows one or the other ancestral line with equal probability, $1/2$. When two lineages meet at a coalescent event in the graph, they necessarily coalesce because $\alpha_{N}\to 1$ under limited outcrossing. Unlinked loci have independent gene genealogies, conditional on the graph. Figure 5 displays simulation results for the SFS for six different values of $\lambda$ ranging from $1000$ to 0. Each panel shows the SFS for five independently generated ancestral graphs assuming a sample of size $n=20$. Given each graph, gene genealogies of $10^{5}$ unlinked loci were simulated and their values of $\tau^{n,r}$ were recorded. Figure 5 plots the averages $\overline{\tau^{n,r}}$ of these branch lengths over the $10^{5}$ gene genealogies, normalized to sum to one, that is divided by $\sum_{r}\overline{\tau^{n,r}}$.
In Figure 5(a), which has $\lambda=1000$, the SFS for every graph is not noticeably different than under the Kingman coalescent (not shown). We note that this is also the expectation for a fixed $\alpha_{N}=\alpha$ ND (97); NWF (25). A small effect of the graph (or pedigree) can be seen even when $\lambda=100$ in Figure 5(b). At the other extreme, when $\lambda=0$ as in Figure 5(f), the entire pedigree of the population is reduced to a single Kingman coalescent tree, and the gene genealogies of every locus have this exact same tree. Here, the SFS reflects the random outcomes, of branch lengths and numbers of descendants of each branch, of an effectively single-locus coalescent process. Something quite similar is seen with $\lambda=0.1$ in Figure 5(e). The middle plots, Figure 5(c) and Figure 5(d), with $\lambda=10$ and $\lambda=1$, occupy a transition zone between the deterministic case ($\lambda\to\infty$) where all graphs have the same SFS and the highly stochastic case ($\lambda\to 0$) where each random graph has a distinctly different SFS.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: Expected site-frequency spectra under the limited-outcrossing model for 55 independently generated ancestral graphs, corresponding to 55 pedigrees, for each of six different values of $\lambda\in\{1000,100,10,1,0.1,0\}$ in decreasing order from panel (a) to (f). SFS on the vertical axis means the relative expected number of polymorphic sites where the mutant allele is found in each count $r\in\{1,2,3,\ldots,19\}$ in a sample of size $n=20$. These were estimated by simulating gene genealogies of $10^{5}$ unlinked loci on each graph/pedigree.
5.3 Diploid Sargasyan-Wakeley model
Consider our diploid Sargasyan-Wakeley model in Example 2.2. In this case (3.6) reduces to
$c_{N}=\mathbb{E}\left[\frac{K}{N}\frac{N-K}{N-1}\frac{1}{N}\,+\,\frac{K(K-1)}{N(N-1)}\,\frac{1}{2P}\right],$ (5.2)
where we noted that $2\frac{K}{N}\frac{N-K}{N-1}$ and $\frac{K(K-1)}{N(N-1)}$ are respectively the probability that exactly one (of the two distinct individuals) is a child and the probability that both individuals are children.
To apply the Theorem 4.7, we need to find assumptions on the joint distribution $\pi_{N}$ of $(K_{N},P_{N})$ under which Assumption 3.6 holds, and for which we can explicitly describe the governing measure $\Xi$. To this end, we write $\pi_{N}$ as a mixture on the unit square $0,1]^{2}$. We assume that there exists a measure $\nu_{N}$ on $0,1]^{2}$ so that
$\pi_{N}:=\int_{[0,1]^{2}}\delta_{\lfloor\bar{K}N\rfloor}\otimes\delta_{\lfloor\bar{P}N\rfloor}d\nu_{N}(\bar{K},\bar{P}).$
Define a family of measure $\mu_{N}$ on $0,1]\times\mathbb{N}$ by
$\mu_{N}(dx,m):=\nu_{N}\left(dx\times\left[\frac{m}{N},\frac{m+1}{N}\right)\right).$
Theorem 5.5.
Suppose, as $N\to\infty$, that Assumptions 3.3 and 3.4 hold, and that $c_{N}^{-1}\mu_{N}$ converges vaguely on $0,1]\times\mathbb{N}$ to a finite measure $\mu$. Define the measure $\Xi^{\prime}(\mu)$ on $\Delta$ by
$\int_{(0,1]\times\mathbb{N}}\frac{x^{2}}{m}\,\delta_{\big(\frac{x}{m},\ldots,\frac{x}{m},0,0,\ldots\big)}\,d\mu(x,m),$
where the dirac mass has $m$ copies of $\frac{x}{m}$ in a row and then all zeroes. Then, where $\Pi$ is an EFC with $c_{k}=2(1-\Xi^{\prime}(\mu)(\Delta))$ and $\mu_{\nu_{\rm Coag}}=2\Xi^{\prime}(\mu)$, we have that $\mathcal{L}^{N,n}$ converges weakly as a random measure to $\mathcal{L}_{\Pi}^{n}$.
Proof.
To prove the claim we proceed by characterizing the ordered offspring distribution $\Phi_{N}$ on $\Delta$ from Lemma Appendix A.1. That is, let $V_{(1)}$, $V_{(2)}$, $\ldots$, $V_{(N)}$ denote the ordered genetic contribution of individuals between time-steps. $\Phi_{N}$ is defined to be the law of the infinite tuple in $\Delta$
$\left(\frac{V_{(1)}}{2N},\frac{V_{(2)}}{2N},\ldots,\frac{V_{(N)}}{2N},0,0,\ldots\right).$
Lemma Appendix A.1 would imply that Assumption 3.6 holds, and that $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent where $\Xi=2\left(\Xi^{\prime}(\mu)+(1-\Xi^{\prime}(\mu))\delta_{\mathbf{0}}\right)$. An application of Theorem 4.7 then gives the result. It suffices, therefore, to show that the conditions of Lemma Appendix A.1 hold.
We will see that, conditional on $(x,m)$ in $(0,1]\times\mathbb{Z}_{+}$, the difference between $\frac{V_{(1)}}{2N}$ and $\frac{V_{(m)}}{2N}$ is negligible as $N\to\infty$. We demonstrate this with the slightly simpler case of putting balls into boxes. The variance of putting $\lfloor xN\rfloor$ balls into $m$ buckets is of order $\frac{\lfloor xN\rfloor}{m}$. When renormalizing by $2N$, the variance is of order $O(\frac{1}{N})$. This shows that the difference between the proportion of the $N$ balls when $\lfloor xN\rfloor$ of the balls are dropped into $P$ boxes for the box with the most and the box with the least vanish as $N\to\infty$. Further, this proportion converges to $\frac{x}{m}$. This is not the exact process for $V_{(i)}$, though the argument carries over in much the same way regardless of which $\alpha$ to which $\alpha_{N}\to\alpha$. However, as $\alpha_{N}\to 1$ by Assumption 3.4 we have that the comparison can be made directly to balls in boxes simply by disregarding the proportion of balls that are the result of outcrossing.
By the above reasoning and the assumption that $c_{N}^{-1}\mu_{N}\to\mu$, Lemma Appendix A.1 yields that Assumption 3.6 holds for $Q_{n}$ the infinitesimal generator of an $n$-$\Xi$-coalescent governed by $\Xi=2(\Xi^{\prime}(\mu)+(1-\Xi^{\prime}(\mu))\delta_{\mathbf{0}})$. This gives the claim.
5.4 Mixed demographies
Many models of interest involve mixed demographic behavior, such as Kingman-like behavior with rare extreme events, as in BBE (13); DFBW (24). Suppose that we have two different parent-child distributions $\pi_{N}^{(1)}$ and $\pi_{N}^{(2)}$ on $N]\times[N]$ for the Sargasyan-Wakeley model of Example 2.2. Let $c_{N}^{(1)}$ and $c_{N}^{(2)}$ denote the one-step transition probabilities for two sample lineages in distinct individuals coalescing in a single time-step. Similarly, we let $d_{N}^{(1)}$ and $d_{N}^{(2)}$ denote the outcrossing probabilities for a single time-step for each of the two demographies. For simplicity, we assume that the selfing probabilities $\alpha_{N}^{(1)}$ and $\alpha_{N}^{(2)}$ for the two demographic histories are identical, i.e. $\alpha_{N}^{(1)}=\alpha_{N}^{(2)}=\alpha_{N}$. For any fixed $\rho>0$ we define the mixed parent-child distribution $\pi_{N}$ by
$\pi_{N}=(1-\rho c_{N}^{(1)})\pi_{N}^{(1)}+\rho c_{N}^{(1)}\pi_{N}^{(2)}.$
Finally, let $\mathcal{A}_{N}^{\rm mixed}$ denote the random pedigree governed by the mixed demography $\pi_{N}$ and $\bar{\chi}^{N,n}$ denote the coalescent with time rescaled by $c_{N}^{(1)}$, as in Equation (3.7). The quenched limit of gene genealogies for the mixed demography is characterized by the following corollary.
Corollary 5.6.
Suppose that, as $N\to\infty$, $c_{N}^{(1)}\to 0$, $\rho c_{N}^{(2)}\to\rho^{\prime}<\infty$, that $\pi_{N}^{(i)}$ satisfies Assumption 3.6 and the measure on $\Delta$ governing the coalescences of the demography of $\pi_{N}^{(i)}$ is $\Xi^{(i)}$ for each $i\in\{1,2\}$, that $\left(c_{N}^{(1)}\right)^{-1}d_{N}^{(1)}\to\lambda<\infty$, and that $d_{N}^{(2)}\to 0$. Then the random measure
$\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm mixed}\right)$
converges weakly in distribution to $\mathcal{L}_{\Pi}^{n}$, where $\Pi$ is an EFC with characteristics $c_{k}=\Xi^{(1)}(\mathbf{0})+\rho^{\prime}\Xi^{(2)}(\mathbf{0})$, $\nu_{Coag}=\Xi^{(1)}+\rho^{\prime}\Xi^{(2)}-c_{k}\delta_{\mathbf{0}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$.
Proof.
As the two demographies satisfy Assumption 3.6 so to does $\pi_{N}$. If $Q_{n}^{(i)}$ is the generator of the demography $\pi_{N}^{(i)}$ then we have that the generator of the mixed demography $Q_{n}^{\rm mixed}$ is
$Q_{n}^{\rm mixed}=Q_{n}^{(1)}+\rho^{\prime}Q_{n}^{(2)}.$
Observe that consistency of generators is closed under linearity, and so $\left(Q_{n}^{\rm mixed}\right)_{n\in\mathbb{N}}$ is a consistent family of generators. Further, the relationship connecting infinitesimal generators on $\mathcal{E}_{\infty}$ and measures on $\Delta$ described in Remark 3.8 is linear. Hence there is a unique measure $\Xi^{\rm mixed}$ associated to the family $\left(Q_{n}^{\rm mixed}\right)$ given by
$\Xi^{\rm mixed}=\Xi^{(1)}+\rho^{\prime}\Xi^{(2)}.$
As $d_{N}^{(1)}(c_{N}^{(1)})^{-1}\to\lambda<\infty$ and $d_{N}^{(2)}\to 0$ the outcrossing probability of the mixed demography $d_{N}^{\rm mixed}$ satisfies $d_{N}^{\rm mixed}\left(c_{N}^{(1)}\right)^{-1}\to\lambda$. The result then follows directly from Theorem 4.7. ∎
As a particular application of Corollary 5.6, we adapt the model of BBE (13). In the model of BBE (13) we have that
$\pi_{N}^{\rm BBE}=\left(1-\frac{\rho}{N(N-1)}\right)\delta_{1}\otimes\delta_{N}+\frac{\rho}{N(N-1)}\delta_{\lfloor\psi N\rfloor}\otimes\delta_{2},$ (5.3)
where $\psi$ is a fixed number in $[0,1]$. That is, with probability $1-\frac{\rho}{N^{2}}$ a single time-step consists of a single reproductive event, a single child born to any one of the $N$ potential parents with probability $\alpha_{N}$ or to any of the $\binom{N}{2}$ pairs of parents with probability $1-\alpha_{N}$; that is, it behaves as a Moran model as in NWF (25). With probability $\frac{\rho}{N^{2}}$, two parents reproduce to form $\lfloor\psi N\rfloor$ of the individuals in the nest time-step (going forward in time.) Let $\mathcal{A}_{N}^{\rm BBE}$ denote the pedigree associated to $\pi_{N}^{\rm BBE}$.
Corollary 5.7.
Suppose that, as $N\to\infty$, we have that $(1-\alpha_{N})N^{-1}\to\lambda<\infty$. Then the quenched law of the time-rescaled coalescent associated to $\pi_{N}^{\rm BBE}$
$\mathbb{P}\left(\tilde{\chi}^{N,n}\left(\lfloor t\left(c_{N}^{(1)}\right)^{-1}\rfloor\right)\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm BBE}\right)$
converges weakly in distribution as a random measure to $\mathcal{L}_{\Pi}^{n}$, where $\Pi$ is the EFC of Example 4.3.
Proof.
We follow the notation introduced for Corollary 5.6, where $\pi_{N}^{(1)}=\delta_{1}\otimes\delta_{2}$ and $\pi_{N}^{(2)}=\delta_{\lfloor\psi N\rfloor}\otimes\delta_{2}$. With this notation we have that
$c_{N}^{(1)}=\frac{1}{N(N-1)}\quad\text{ and }c_{N}^{(2)}=\frac{\psi^{2}}{2}+o(1).$
Therefore we may write Equation (5.3) as
$\pi_{N}^{\rm BBE}=(1-\rho c_{N}^{(1)})\pi_{N}^{(1)}+\rho c_{N}^{(1)}\pi_{N}^{(2)},$
and we see that $\rho c_{N}^{(2)}\to\rho\frac{\psi^{2}}{2}$
Observe now that the probability of three sample lineages in three distinct individuals coalescing in a single time-step for $\pi_{N}^{(1)}$ is zero. One readily checks that the assumptions of Theorem 5.2 are satisfied and hence $\Xi^{(1)}=2\delta_{\mathbf{0}}$.
We now calculate the transition-rate matrix $Q_{l}^{(2)}=\left(q^{l,2}_{\xi\eta}\right)_{\xi,\eta\in\mathcal{E}_{l}}$ for $\pi_{N}^{(2)}$. It suffices to do so for $q_{\xi_{0}^{l}\eta}^{l}$ for a fixed $l\in\mathbb{Z}_{+}$ by exchangeability. Observe firstly that if $\eta$ consists of at least three non-singleton blocks, then we immediately have that $q_{\xi_{0}^{l}\eta}^{l}=0$. This is simply because their are precisely two parents under $\pi_{N}^{(2)}$.
Suppose now that $\eta$ consists of $b$ blocks $C_{,}C_{2},...,C_{b}$ of sizes $k_{1},k_{2},\ldots,k_{b}$, at most two of which are non-singletons. Where $s$ is the number of singletons, we write that $\eta$ is of type $b;k_{1},\ldots,k_{r};s$, as in Sch (00). We only need to consider the cases now where $s\geq b-2$. Observe that the probability that any of the $l$ individuals containing a sample lineage is one of the two parents is $o(1)$, which will be negligible for $\pi_{N}^{(2)}$. We proceed now assuming that $k_{1}\geq 2$ and that $k_{1}\geq k_{2}\geq\ldots\geq k_{b}$ without loss of generality.
For any $j$ of the $l$ sample lineages, the probability that they are the child of one of the two parents is $\psi^{j}(1-\psi)^{l-j}+o(1)$. If $k_{2}\geq 2$, then there is only one way for $\eta$ to appear, by $k_{1}$ sample lineages falling into one parent, $k_{2}$ in the other, and the remaining $l-k_{1}-k_{2}$ remaining sample lineages not being children. Therefore
$q^{l,2}_{\xi_{0}^{l}\eta}=2^{1-k_{1}-k_{2}}\psi^{k_{1}+k_{2}}(1-\psi)^{l-k_{1}-k_{2}}+o(1).$ (5.4)
The factor of $2$ comes from the $k_{1}+k_{2}$ sample lineages sorting themselves into the right parents, where the order of the parents does not matter.
If $k_{2}=1$, then $k_{1}$ sample lineages fall into one of the two parents. It is possible that exactly one or 0 of the remaining $l-k_{1}$ sample lineages fall into the other parent. Therefore
$q^{l,2}_{\xi_{0}^{l}\eta}=2\left(\frac{\psi}{2}\right)^{k_{1}}\left(\frac{\psi}{2}(1-\psi)^{l-k_{1}-1}+(1-\psi)^{l-k_{1}}\right)+o(1)=2^{1-k_{1}}(1-\psi)^{l-k_{1}}\frac{2-\psi}{2-2\psi}+o(1).$ (5.5)
Note that the first term in the summand comes from one of the $l-k_{1}$ remaining sample lineages falling into one of the two parents, while the second term comes from calculating that none of the remaining sample lineages do.
Finally, we calculate $q^{l,2}_{\xi_{0}^{l}\xi_{0}^{l}}$. There can be zero, one, or two of the $l$ sample lineages that enter the two parents. Zero lineages entering the two parents occurs with probability $(1-\psi)^{l}+o(1)$. There are $\binom{l}{1}$ ways for $1$ sample lineage to be in one of the two parents, and so this occurs with probability $\binom{l}{1}\psi(1-\psi)^{l-1}+o(1)$. There are $\binom{l}{2}$ ways for $2$ sample lineages to be in the two parents, and in this case they do not coalesce together with probability $\frac{1}{2}$. Therefore
$q^{l,2}_{\xi_{0}^{l}\xi_{0}^{l}}=(1-\psi)^{l}+l(1-\psi)^{l-1}\psi+\frac{l(l-1)}{4}(1-\psi)^{l-2}\psi^{2}+o(1).$ (5.6)
Combining Equations (5.4), (5.5), and (5.6) shows that $Q_{n}^{(2)}$ corresponds exactly to the measure $\Xi^{(2)}=\delta_{\left(\frac{\psi}{2},\frac{\psi}{2},0,\ldots\right)}$. The result then follows from Corollary 5.6. ∎
Having stated our main result and its applications, we organize the rest of the paper as follows: In Section 6 we provide a simple criterion by which weak convergence of random measures over the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ converges when $E$ is taken to be a locally compact Polish space. In Section 7 we show that that the subgraph of the pedigree consisting of possible lineage trajectories converges to a $Q$-$\lambda$ graph, which is then realized via a coupling argument as a subgraph of an EFC. With the characterization of weak convergence of random measures in hand, the quenched convergence of coalescing random walks on the pedigree is shown by a continuity argument.
6 Characterization of weak convergence in distribution of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$
Possible references for what follows are BK10a ; BK10b ; Led (16); Mit (83).
To prove Theorem 4.7 we require a characterization of weak convergence in distribution of random measures. Following (Kal, 17, Theorem 4.19, pp. 111, 126), we provide the following definition of weak convergence in distribution for random probability measures.
Definition 6.1.
Let $\mu$ and $(\mu_{N})_{N\in\mathbb{N}}$ denote random probability measures on a Polish space $S$. Then $\mu_{N}$ converges weakly in distribution to $\mu$ if for any continuous $f$ of bounded support that
$\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f).$
When this convergence holds we write $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mu$.
We provide in this section sufficient criteria to characterize weak convergence in distribution when the Polish space $S$ in question is precisely $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ with the $J_{1}$ topology, where $E$ is a locally compact Polish space. The basic strategy is to construct a suitable dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ and then to devise criteria by which evaluations our random measures against this suitable class of test functions converge.
6.1 A dense family in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$
Suppose that $E$ is a locally compact Polish space. Then $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ is a Polish space under the $J_{1}$ topology (EK, 09, Theorem 5.6, p.121). We shall describe a dense subset of $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ in Lemma 6.4 below when $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ is given the topology of uniform convergence on compacta. To this end, we first describe an explicit continuous function on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
Fix $\varphi\in C_{c}\left(\mathbb{R}_{+}\times E\right)$ and $a\in\mathbb{R}$, and define the function $I(\varphi,a):\,\mathcal{D}\left(\mathbb{R}_{+},E\right)\to\mathbb{R}$ by
$I(\varphi,a)(x):=a+\int_{0}^{\infty}\varphi(s,x(s))ds.$
Here, for any topological space $X$ we take $C_{c}(X)$ to denote the continuous functions from $X$ into $\mathbb{R}$ with compact support.
Lemma 6.2.
For any $\varphi\in C_{c}\left(\mathbb{R}_{+}\times E\right)$ and $a\in\mathbb{R}$, $I(\varphi,a)$ is a bounded and continuous function on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
Proof.
We need to show that for any sequence $x_{N}$ in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ converging to a path $x$ therein that $I(\varphi,a)(x_{N})$ converges to $I(\varphi,a)(x)$.
As $\varphi$ has compact support, there exists a finite $T$ in $\mathbb{R}_{+}$ such that $supp(\varphi)\subset[0,T]\times E$. Therefore
$I(\varphi,a)(x_{N})=a+\int_{0}^{T}\varphi(s,x_{N}(s))ds.$
[…]
We now define the collection $\mathcal{I}:=\{I(\varphi,a):\,\varphi\in C_{c}\left(\mathbb{R}_{+}\times E\right),\,a\in\mathbb{R}\}$. Note that $I(\varphi,a)\subset C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ by Lemma 6.2.
To apply a Stone-Weierstrass, we begin by showing that $\mathcal{I}$ separates points in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. That is, for any $x\neq y$ there exists $\varphi,a$ such that $I(\varphi,a)(x)\neq I(\varphi,a)(y)$.
Lemma 6.3.
For any locally compact Polish space $E$, $\mathcal{I}$ separates points in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
Proof.
If $x\neq y$, then there is a continuity point $t_{0}\in\mathbb{R}_{+}$ of both $x$ and $y$ such that either $x(t_{0})\neq y(t_{0})$. Because $E$ is locally compact and Hausdorff, there exist disjoint open neighborhoods with compact closure $U$ and $V$ of $x(t_{0})$ and $y(t_{0})$, respectively. Choose $\theta$ in $C_{c}(E,[0,1])$ with $\theta_{|U}\equiv 1$ and $\theta_{|V}\equiv 0$, which exists by Urysohn’s lemma.
As $t_{0}$ is a continuity point, there is a $\delta>0$ such that $x(s)\in U$ and $y(s)\in V$ for all $|s-t_{0}|<\delta$. Let $\psi\in C_{c}(\mathbb{R}_{+},[0,1])$ satisfy $\psi_{|[t_{0}-\frac{1}{2}\delta,t_{0}+\frac{1}{2}\delta]}\equiv 1$ and $supp(\psi)\subset[t_{0}-\delta,t_{0}+\delta]$.
Define $\varphi$ by
$\varphi(s,e):=\psi(s)\theta(e).$
$\varphi$ clearly belongs to $C_{c}(\mathbb{R}_{+}\times E)$. Further
$I(\varphi,0)(x)>0=I(\varphi,0)(y),$
so $I(\varphi,0)$ separates $x$ and $y$. Thus $\mathcal{I}$ separates points of $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. ∎
We now define $\bar{\mathcal{I}}$ to be the collection of all finite sums of finite products of elements of $\mathcal{I}$. That is, each element $I$ in $\bar{\mathcal{I}}$ may be written as
$\sum_{i=1}^{n}\prod_{j=1}^{k_{i}}I_{ij},$
where each $I_{ij}$ belongs to $\mathcal{I}$ and $k_{i},n\in\mathbb{Z}_{+}$. An extension of this class of continuous, bounded integral functionals will, by an application of Stone-Weierstrass Ok (25)[Chapter 7, p. 11], be dense in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
Lemma 6.4.
Suppose that $E$ is locally compact and Polish. Then $\bar{\mathcal{I}}$ is dense in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ in the topology of uniform convergence on compacta.
Proof.
To show that $\bar{\mathcal{I}}$ is dense in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ in the topology of uniform convergence on compacta, we need to show that for any function $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$, any $\varepsilon>0$, and any compact set $K$ there exists a function $I$ in $\bar{\mathcal{I}}$ such that
$\sup_{e\in K}|f(e)-I(e)|<\varepsilon.$
We begin by observing that $\bar{\mathcal{I}}$ contains constants; take $\varphi\equiv 0$ and $a$ to be free. $\bar{\mathcal{I}}$ also is closed under addition, multiplication, and scalar multiplication, which is immediate. By Lemma 6.3 $\mathcal{I}$, and so $\bar{\mathcal{I}}$, separates points.
Fix any compact set $K\subset\mathcal{D}\left(\mathbb{R}_{+},E\right)$ and any $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$. The restriction of elements of $\bar{\mathcal{I}}$ to $K$ yields an algebra $\bar{\mathcal{I}}(K)$ of continuous functions from $K$ to $\mathbb{R}$ that contains the constants and separates points. By the Stone-Weierstrass theorem for compact Hausdorff spaces Ok (25)[Chapter 7, p. 11], the uniform closure of $\bar{\mathcal{I}}(K)$ is all of $C(K)$. This gives the claim.
6.2 A characterization of weak convergence in distribution
We show by means of a density argument that it suffices to check, for weak convergence in distribution of random measures, joint convergence in distribution when testing against integral functionals, as described by the following Theorem.
Theorem 6.5.
Let $\mu_{N}$ denote a sequence of random measures in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ for a locally compact, Polish space $E$. Then $\mu_{N}$ converges weakly in distribution to $\mu$ if and only if for any finite collection $\{\varphi_{i}\}_{i=1}^{k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times E\right)$ that
$\left(\mu_{N}\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\left(\mu\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}.$ (6.1)
Proof.
From Definition 6.1 it suffices to show that for any $f$ in $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ that
$\mu_{N}(f)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mu(f).$ (6.2)
By Lemma 6.4 it suffices, without loss of generality, to demonstrate (6.2) for $f$ in $\bar{\mathcal{I}}$. For any element $f$ in $\bar{\mathcal{I}}$ there is an array of elements $\left(I_{ij}\right)_{i\in[n],j\in[k_{i}]}$ of $\mathcal{I}$ such that
$f=\sum_{i=1}^{n}\prod_{j=1}^{k_{i}}I_{ij}.$
Linearity of expectation yields
$\mu_{N}(f)=\sum_{i=1}^{n}\mu_{N}\left(\prod_{j=1}^{k_{i}}I_{ij}\right).$ (6.3)
By assumption, $\mu_{N}(\left(I_{ij}\right)_{i\in[n],j\in[k_{i}]})$ converges in distribution to $\mu(\left(I_{ij}\right)_{i\in[n],j\in[k_{i}]})$. As polynomials are continuous and continuity preserves convergence in distribution, (6.3) converges in distribution to
$\sum_{i=1}^{n}\mu\left(\prod_{j=1}^{k_{i}}I_{ij}\right)=\mu(f),$
which gives the claim.
We now prove the necessity of Equation (6.1). By Lemma 6.2 the integral functionals are all continuous. The continuous mapping theorem then gives the claim.
The characterization of weak convergence in distributions of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ for locally compact and Polish $E$ given by Theorem 6.5 allows one to give classical-like characterizations of this form of weak convergence, as typified by the following following.
Definition 6.6 (Weak convergence in finite-dimensional distribution).
Let $\mu$, $(\mu_{N})_{N\in\mathbb{N}}$ denote random probability measures over the space $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. The sequence $\mu_{N}$ converges weakly in finite-dimensional distribution to $\mu$ if for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$ the pushforward by the evaluation map
$\pi_{\vec{t}}:\mathcal{D}\left(\mathbb{R}_{+},E\right)$ $\to E^{k}$
$x$ $\mapsto\left(x(t_{i})\right)_{1\leq i\leq k}$
converges weakly in distribution as random measures. That is,
$\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu.$
We know define an analogue of the compact containment condition for random measures on Skorokhod space (EK, 09, p. 129).
Definition 6.7 (Compact containment condition for random measures).
Let $\mu_{N}$ denote a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. Then $\mu_{N}$ satisfies the compact containment condition if, for every $\varepsilon>0$ and $T>0$ that there is a compact set $K(\varepsilon,T)\subset E$ such that
$\mathbb{P}\left(\limsup_{N\to\infty}\mu_{N}\left(x([0,T])\subset K(\varepsilon,T)\right)\geq 1-\varepsilon\right)=1.$
We obtain the following characterization of weak convergence of random measures in $\mathcal{D}\left(\mathbb{R}_{+},E\right)$.
Theorem 6.8.
Suppose $\mu$ is a random measure on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$, where $E$ is a locally compact Polish space. Suppose now that $\mu_{N}$ is a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ such that
  • for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$ we have that $\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{t}\mu$,
  • and $\mu_{N}$ satisfies the compact containment condition for random measures from Definition 6.7,
Then $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mu$.
Proof.
By Theorem 6.5 it suffices to show that, for any finite collection $\left(\varphi_{i}\right)_{1\leq i\leq k}$ of elements of $C_{c}\left(\mathbb{R}_{+},E\right)$ that
$\left(\mu_{N}\left(\int_{0}^{\infty}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\left(\mu\left(\int_{0}^{\infty}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}.$
As each $\varphi_{i}$ is compactly supported, there exists $T>0$ such that $\cup_{i}supp\left(\varphi_{i}\right)\subset[0,T]\times E$. By the compact containment condition, for any $\eta,\rho>0$ there exists $K(\eta,T)$ a compact set in $E$ such that
$\mathbb{P}\left(\mu_{N}\left(x([0,T])\subset K(\eta,T)\right)\geq 1-\eta\right)\geq 1-\rho$ (6.4)
for all sufficiently large $N$.
Since $\varphi_{i}$ is continuous on $0,T]\times K(\eta,T)$ for any compact $K\subset E$, it is uniformly continuous there. Fix now $\delta>0$. Choose a fine partition $0=t_{0}<t_{1}<\ldots<t_{r}=T$ so that $\min_{i}\{t_{i+1}-t_{i}\}<\delta$. By uniform continuity, for any $\varepsilon>0$ we can choose $\delta$ so that
$|\varphi_{i}(s,e)-\varphi_{i}(t,e)|<\varepsilon$
for any $|s-t|<\delta$.
For any path $x$ for which $x(s)\in K(\eta,T)$ for all $0\leq s\leq T$ we have that
$\left|\int_{0}^{T}\varphi_{i}\left(s,x(s)\right)-\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right|\leq T\varepsilon.$ (6.5)
For all other $x$ we have that
[…]
Combining (6.5), (6.6) yields
[…]
By Markov’s inequality and (6.4) we therefore have, for any fixed $\theta>0$, that
[…]
As $\eta,\rho,\varepsilon,\theta$ are arbitrary, we therefore have that
$\mu_{N}\left(\left(\left|\int_{0}^{T}\varphi_{i}\left(s,x(s)\right)-\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right|\right)_{1\leq i\leq k}\in\cdot\right)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\delta_{0}$
as $N$ goes to infinity.
As $\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu$, the vector of Riemann sums satisfies
$\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\left(\left(\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right)_{1\leq i\leq k}\in\cdot\right)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu\left(\left(\sum_{j=0}^{r-1}\left(t_{j+1}-t_{j}\right)\varphi_{i}(t_{j},x(t_{j}))\right)_{1\leq i\leq k}\in\cdot\right).$
This gives the claim.
When $E$ is compact, we immediately have the following corollary.
Corollary 6.9.
Suppose that $E$ is a compact metric space and $\mu$ is a random measure on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$. Let $\mu_{N}$ denote a sequence of random measures on $\mathcal{D}\left(R_{+},E\right)$ such that, for any finite collection of times $\left(t_{i}\right)_{1\leq i\leq k}$ such that $\left(\pi_{\vec{t}}\right)_{*}\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\left(\pi_{\vec{t}}\right)_{*}\mu$. Then $\mu_{N}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mu$.
Remark 6.10.
Corollary 6.9 allows us to strengthen some already existing results, including (ABFW, 25, Theorem 3.8). Theorem 3.8 demonstrates quenched convergence in finite-dimensional distribution to a time-inhomogeneous coalescent. By the corollary this is enough to show that the quenched convergence in laws follows.
7 Proof of Theorem 4.7
We will relate subgraphs of an EFC process $\Pi$ to what we call a $Q$-$\lambda$ graph. Here $Q$ corresponds to the generator of a $\Xi$-coalescent, $\lambda$ refers to the fragmentation rate of the particles. In Section 7.1 we describe $Q$-$\lambda$ graphs as particle systems where an initial sample of size $n$, each particle fragments into two particles independently at rate $\lambda$, and where particles coalesce as a $\Xi$-coalescent. In Section 7.2 we show that the subgraph of the pedigree given by following all possible trajectories of a sample of $n$ lineages converges, with the $c_{N}^{-1}$ time-rescaling to a $Q$-$\lambda$ graph. In Section 7.4 we show that the coalescent of coalescing random walks on the discrete-time ancestral graph converges to that of a $Q$-$\lambda$ graph. By realizing the $Q$-$\lambda$ graph as a subgraph of an EFC process via a coupling argument, we finish the proof of Theorem 4.7.
7.1 Construction of the $Q$-$\lambda$ graph
Fix $\lambda\in\mathbb{R}_{+}$ and a sequence of consistent matrices $\left(Q_{n}\right)_{n\in\mathbb{N}}$, and write $Q_{n}=\left(q_{\xi\eta}^{n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$. We can assume they are obtained from Assumptions 3.4 and 3.6 respectively. We shall define a stochastic process, called a a $Q$-$\lambda$ ancestral graph, with state space
$\Theta:=\mathbb{Z}_{+}\times\mathbb{Z}_{+}\times\mathcal{E},$
where $\mathcal{E}=\cup_{n\geq 2}\mathcal{E}_{n}$. For each $(l,m,\xi)\in\Theta$, the integer $l$ will track the current number of lineages, $m$ the label of the particle that most recently fragmented, and $\xi$ the structure of the most recent coalescence event.
This process is a random graph in which each node in the graph fragments independently at rate $\lambda$, and where any subset of $r$ nodes, represented by the singleton $\xi_{0}^{r}=\{\{i\}\}_{i=1}^{r}$, coagulate according to $\eta\in\mathcal{E}_{r}$ with rate $q_{\xi_{0}^{r}\eta}^{r}$. We provide the formal definition below.
Definition 7.1.
A $Q$-$\lambda$ ancestral graph is a continuous-time Markov process with sample paths in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ and infinitesimal generator $\mathscr{L}$ acting on $f\in C_{b}(\Theta)$ by
$\mathscr{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}q_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.1)
For $n\geq 1$, we denote by $G_{Q,\lambda}^{n}=\big(G_{Q,\lambda}^{n}(t)\big)_{t\in\mathbb{R}_{+}}$ a $Q$-$\lambda$ ancestral graph starting with $n$ nodes, i.e. when $G_{Q,\lambda}^{n}(0)=(n,m,\xi)$ for some $(m,\xi)\in\mathbb{Z}_{+}\times\cup_{n\geq 2}\mathcal{E}_{n}$.
Existence of $G_{Q,\lambda}^{n}$ holds because fragmentation rate is linear in the first coordinate and does not depend on the the other two coordinates. This can be verified by, for instance, (EK, 09, Proposition 2.9 in Chap. 4).
We describe now $n$ independent random walks $\{x_{i}\}_{1\leq i\leq n}$ on the ancestral graph $G_{Q,\lambda}^{n}$, which we view as $\mathbb{Z}_{+}$-valued processes. The joint process $\left(G_{Q,\lambda}^{n},x_{1},\ldots,x_{n}\right)$ may be described by a generator $\tilde{\mathscr{L}}$ defined on $C_{b}\left(\Theta\times\mathbb{Z}_{+}^{n}\right)$. Define a vector $\sigma:\mathbb{Z}_{+}^{n}\times\mathbb{Z}_{+}\to\{0,1\}^{n}$ by
$\sigma(x,j)_{i}:=\mathbbm{1}_{\{x_{i}=j\}}.$
The vector $\sigma(x,j)$ in the $i$th position is $1$ if $x_{i}=j$ and 0 otherwise. This gives a compact representation of all the walks that currently are in position $j$ on the graph.
Define another vector $\rho:\mathbb{Z}_{+}^{n}\times\mathcal{E}\to\mathbb{Z}_{+}^{n}$ where $\rho(x,\xi)_{i}$ is equal to the index of the block to which the $i$th particle belongs. To make this canonical, we may suppose that $\xi=\{C_{1},C_{2},\ldots,C_{b}\}$ is ordered by its least element. That is, for each $r<s$ we have that
$\inf C_{r}<\inf C_{s}.$
$\rho(x,\xi)_{i}$ is therefore equal to the index of the unique block $C_{\rho(x,\xi)_{i}}=[x_{i}]_{\xi}$ of $\xi$, where $x_{i}]_{\xi}$ is the equivalence class of $x_{i}$ with respect to $\xi$.
Definition 7.2 (Coalescing random walks on the ancestral graph).
The $n$ $\mathbb{Z}_{+}$-valued random walks $\{x_{i}\}_{1\leq i\leq n}$ on the $Q$-$\lambda$ ancestral graph $G_{Q,\lambda}^{n}$ have their joint distribution with $G_{Q,\lambda}^{n}$ described by the generator $\tilde{\mathscr{L}}$. They satisfy that
  1. 1.
    $x_{i}(0)=i$ for all $1\leq i\leq n$,
  2. 2.
    $x_{i}(t)\leq l(t)$ for all $1\leq i\leq n$ and all $t\in\mathbb{R}_{+}$,
  3. 3.
    at any fragmentation event, each of the $x_{i}$ will become $l+1$ or remain unchanged with equal (i.e. $\frac{1}{2}$) probability,
  4. 4.
    once any two of the $n$ processes agree, they remain identical for all time thereafter.
This system of coalescing random walks on the $Q$-$\lambda$ graph gives rise to a unique (in law) process with sample paths in $\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$, which we denote by $\chi^{n}_{G}$, when we keep track of the indices of these random walks.
Definition 7.3.
We denote the law of this process $\chi^{n}_{G}$, conditional on $G_{Q,\lambda}^{n}$, by
$\mathcal{L}_{Q,\lambda}^{n}:=\mathbb{P}\left(\chi^{n}_{G}\in\cdot\,\mid\,G_{Q,\lambda}^{n}\right).$
7.2 Convergence of the discrete-time ancestral graph
We begin by giving a construction of the discrete-time ancestral graph $G^{N,n}=\left(l^{N},m^{N},\xi^{N}\right)$ as a discrete-time process taking values in $\Theta$.
We will define $G^{N,n}$ in terms of another process, $\tilde{G}^{N,n}$ which is the subgraph of the population pedigree $\mathcal{G}_{N}$ given by all possible trajectories of the sample lineages $\left(\hat{X}_{i}^{N}\right)_{1\leq i\leq n}$, tracing all possible ancestral individuals of the sample as vertices of this subgraph. Let $\tilde{V}_{k}^{N}\subset[N]$ be the set of all possible ancestral individuals of the sample at time-step $k$ in the past.
That is, $\tilde{G}^{N,n}$ consists of a sequence of edges $\left(\tilde{E}_{k}^{N}\right)$ where the edges from $\tilde{V}_{k}^{N}$ to $\tilde{V}_{k+1}^{N}$ consist of those edges from individual to itself for persisting in that time-step (i.e. not dying) and also edges from parents to children for reproductive events in this time-step. The edge set $\tilde{E}_{k}^{N}$ is merely the edges of the pedigree connecting those elements of $\tilde{V}_{k}^{N}$.
We recall that $n$ is the sample size, and assume the initial conditions is $G^{N,n}(0)=\left(n,1,\xi_{0}^{n}\right)$. For $k\geq 1$, we define $G^{N,n}(k)=\left(l^{N}(k),m^{N}(k),\xi^{N}(k)\right)$ as follows:
  • $l^{N}(k):=|\tilde{V}_{k}^{N}|$ is the number of distinct nodes in the subgraph $\tilde{G}^{N,n}$ at the $k$th time-step,
  • $m^{N}(k)$ is the index of the node in $\tilde{G}$ the subject of most recent outcrossing, i.e.
[…]
    Remark 7.4.
    Note that while it is possible in our model that more than one outcrossing occurs in each time-step, this will occur with negligible probability over the lifespan of the graph. This is shown in Lemma 7.6.
  • Label the vertices $\tilde{V}_{k}^{N}=\{\tilde{v}_{i}\}_{1\leq i\leq l^{N}(k)}$ in order. We then define $\xi^{N}(k)$ as the unique partition in $\mathcal{E}_{l^{N}(k)}$ defined by the transitive closure of the binary relation $R^{N}(k)$
[…]
Figure 6 shows a realization of the ancestral graph $G^{N,n}$ for a sample of size $n=2$.
Figure 6: Here we see the discrete-time ancestral graph $G^{N,2}$ for a sample of size $n=2$. Between the 0-th and the 11-st time-step $m$ changes because the second node, read left to right, fragments. Between the third and fourth time-step the second and third nodes coalesce, while the remaining three nodes are uninvolved in the coalescence.
We now recall the definition of the haploid map $F_{\rm hap}$ of an element of $\mathcal{S}_{n}$, defined in (3.10). The haploid map $F_{\rm hap}$ maps $\mathcal{S}_{n}$ into $\mathcal{E}_{n}$ by collapsing all elements of the form $\{A,B\}$ into $A\cup B$. Formally, for
$\xi=\{(C_{1},C_{2}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\}$
we define
$F_{\rm hap}\left(\xi\right):=\{C_{1}\cup C_{2},\ldots,C_{2x-1}\cup C_{2x},C_{2x+1},\ldots,C_{b}\}.$
Remark 7.5.
Observe that if all of the reproductive events are via selfing then, then $R^{N}(k)=\xi^{N}(k)$. Notice that this is not the case with probability $O(1-\alpha_{N})$. Further, we can compare the probability of this configuration of occurring with the transition matrix $Q_{l}^{N}.$ Indeed, for each $\eta\in\mathcal{E}_{n}$ the rate at which a coalescence with the structure $\eta$ occurs when there are $l$ extant sample lineages is precisely the sum of all the rates at which $l$ sample lineages in $l$ distinct individuals overlap into individuals with the structure $\eta$. That is, the rate at which $l$ sample lineages coagulate like $\eta$ is precisely the sum of all the rates at which the $l$ sample lineages in $l$ distinct individuals enter a state whose image under the haploid map is $\eta$, i.e. $q^{N,l}_{\xi_{0}^{l}\eta}$.
Assume that the demography of the population in the general model of Section 3.6. We let $Q^{N}=\left(Q^{N}_{n}\right)_{n\in\mathbb{Z}_{+}}$ denote the finite $N$ generator described therein, i.e.
$Q^{N}_{n}=\left(q_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}=c_{N}^{-1}\left(\mathit{p}_{\xi\eta}^{N,n}-\delta_{\xi\eta}\right)_{\xi,\eta\in\mathcal{E}_{n}},$ (7.2)
where $\delta_{\xi\eta}$ is the Kronecker delta function. Recall too that the infinite population limit generator $Q_{n}=\left(q_{\xi\eta}^{n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$. We calculate the one-step transition probabilities for $G^{N,n}$ explicitly in terms of $Q^{N}_{n}$ in the following lemma.
Lemma 7.6.
Suppose that Assumption 3.4 holds. Then the one-step transition probabilities of the Markov chain $\{G^{N,n}(k)\}_{k\in\mathbb{Z}_{+}}$ satisfy the following asymptotic property: given that $G^{N,n}(k)=(l,m,\xi)\in\Theta$, it holds that
[…]
for each $(m^{\prime},l^{\prime},l^{\prime\prime},\eta)$ such that $m^{\prime}\neq m$ and $l^{\prime}\neq l+1$, $l^{\prime\prime}\neq|\eta|$, and $\eta\in\mathcal{E}_{l}$. In the above, the error terms $O(c_{N}),o(c_{N}),O(d_{N}c_{N})$ are all uniform for $(l,m,\xi)\in\Theta$.
Remark 7.7.
Note that a sequence $\{a_{N}\}$ belongs to $O(b_{N})$ if there exists a constant $C$ such that $a_{N}\leq Cb_{N}$ for all $N\geq 1$. A sequence $\{a_{N}\}$ belongs to $o(b_{N})$ if $\lim_{N\to\infty}\frac{a_{N}}{b_{N}}=0$. Note that $O(b_{N})$ and $o(b_{N})$ consist of collections of real-valued sequences and so set-valued inclusions between $O(a_{N})$ and $o(b_{N})$ make sense. We write $O(b_{N})\subset o(c_{N})$ if $\lim_{N\to\infty}\frac{a_{N}}{c_{N}}=0$ for any sequence $\{a_{N}\}\in O(b_{N})$.
Proof.
Observe that the probability of either an outcrossing or a coalescence of the $l$ nodes is bounded above by $2c_{N}\binom{l}{2}+ld_{N}$. By Assumption 3.4, the probability of neither a coalescence nor an outcrossing is $1-l^{2}O(c_{N})$. This gives the first line.
Now we show that the probability that two outcrossings occur in the same time-step occurs with negligible probability, implying line 3. Indeed, the probability that two outcrossings occur is bounded by the probability that any of the $\binom{l}{2}$ possible pairs of particles are both children and both are the result of outcrossing. The probability that any such pair consists of two children is $\mathbb{E}\left[\frac{K_{N}(K_{N}-1)}{N(N-1)}\right]$. The probability that they both are the result of outcrossing is $(1-\alpha_{N})^{2}$. Therefore the probability of at least two outcrossings is at most
$\binom{l}{2}(1-\alpha_{N})^{2}\mathbb{E}\left[\frac{K_{N}(K_{N}-1)}{N(N-1)}\right]\leq\binom{l}{2}d_{N}(1-\alpha_{N}).$
This, together with the fact that $\alpha_{N}\to 1$ (by Assumption 3.4), implies that the probability of at least two outcrossings is bounded by $\binom{l}{2}o(d_{N})$. This gives the third line.
For any $\eta\in\mathcal{E}_{n}$ not equal to $\xi_{0}^{l}$, the probability in one time-step that there is any coalescence of the $l$ nodes in the graph with structure $\eta$ and without outcrossing is
$p_{\xi_{0}^{l}\eta}^{N,l}\alpha_{N}^{l}=c_{N}q_{\xi_{0}^{l}\eta}^{N,l}\alpha_{N}^{l},$
which gives line $5$.
We now show that an outcrossing occurs during a coalescence with negligible probability, implying the fourth line. As $\xi_{0}^{l}$ consists of particles in $l$ distinct individuals, the coalescence rate is independent to the selfing rate. In particular, the probability that we have a coalescence with structure $\eta$ and an outcrossing is therefore $O((1-\alpha_{N})q^{N,l}_{\xi\eta})\subset o(c_{N})$. This gives the fourth line. ∎
Equipped with Lemma 7.6 giving the transition rates of the discrete-time ancestral graph $G^{N,n}$, we are prepared to demonstrated convergence of the sped-up ancestral graph to a $Q$-$\lambda$ graph.
Lemma 7.8.
Suppose that, as $N\to\infty$, Assumptions 3.3, 3.4, and 3.6 hold. Then the sped-up discrete-time ancestral graph $\bar{G}^{N}=\left(G^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)\right)_{t\in\mathbb{R}_{+}}$ converges weakly in the $J_{1}$ topology in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to a $Q$-$\lambda$ ancestral graph $G^{n}_{Q,\lambda}$ with respect to the measures $\mathbb{P}_{\xi_{0}^{n}}^{N}$, where $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is as in Assumption 3.6.
Proof.
We proceed as in the proof of NWF (25)[Lemma 7.6].
Let $T^{(N)}$ denote the linear operator on the space $C_{b}\left(\Theta\right)$ of bounded continuous functions on $\Theta$ defined by $T^{(N)}f(l,m,\xi)=\mathbb{E}\left[f\left(G^{N}(1)\right)\mid G^{N,n}(0)=(l,m,\xi)\right]$. The generator $\mathscr{L}^{N}$ of the discrete-time process $G^{N,n}$ is given by
$\mathscr{L}^{N}f(l,m,\xi)=\left(T^{(N)}-I\right)f(l,m,\xi),$
which can be calculated to sufficient accuracy as in Equation 7.3. Let $\mathscr{L}$ be the infinitesimal generator of the continuous-time process $G_{Q,\lambda}^{n}:=\left(G_{Q,\lambda}^{n}(t)\right)_{t\in\mathbb{R}_{+}}$. That is, by Equation (7.1),
$\mathscr{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}\tilde{q}_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.4)
It follows from Equation (7.3) that for all $f\in C_{b}(\Theta)$ with finite support that,
$\sup_{(l,m,\xi)\in\Theta}\left|c_{N}^{-1}\mathscr{L}^{N}f(l,m,\xi)-\mathscr{L}f(l,m,\xi)\right|\to 0\text{ as $N\to\infty$.}$
Let $\{T(t)\}_{t\in\mathbb{R}_{+}}$ be the semigroup on $C_{b}(\Theta)$ of $G_{Q,\lambda}^{n}$. By EK (09)[Theorem 6.6, Chapter 1] and (7.4), it holds that
$\lim_{N\to\infty}\sup_{0\leq t\leq T}\sup_{(l,m,\xi)\in\Theta}\left|\left(T^{(N)}\right)^{\lfloor tc_{N}^{-1}\rfloor}f(l,m,\xi)-T(t)f(l,m,\xi)\right|=0$ (7.5)
for all $f$ in the domain of $\mathscr{L}$
We now proceed to demonstrate tightness. Observe that, for any $n$ that the set $K_{n}:=[n]\times[n]\times\cup_{2\leq k\leq n}\mathcal{E}_{k}$ is compact. To establish the compact containment condition EK (09)[(7.9), p.129] holds if for any $\varepsilon\in(0,1)$ and $T\in(0,\infty)$, there exists $M(\varepsilon,T)\in\mathbb{Z}_{+}$ such that
$\limsup_{N\to\infty}\mathbb{P}_{\xi_{0}^{n}}^{N}\left(G^{N}(t)\in K_{M(\varepsilon,T)}\,\,\forall 0\leq t\leq T\right)\leq\varepsilon.$ (7.6)
Notice that for all time $t$ we have that
$|\xi^{N}(t)|,|m^{N}(t)|\leq\sup_{0\leq s\leq t}l^{N}(t).$
Therefore, to demonstrate (7.6), it suffices to show that
$\limsup_{N\to\infty}\mathbb{P}_{\xi_{0}^{n}}^{N}\left(\sup_{0\leq t\leq T}l^{N}(t)\leq M(\varepsilon,T)\right)\leq\varepsilon.$ (7.7)
Let $\mathcal{Q}$ denote the generator of the cross-section size $L=\left(L_{t}\right)_{t\in\mathbb{R}_{+}}$ for the limiting graph. Then for any $f$ in $C_{b}(\mathbb{Z}_{+})$ we have that
$\mathcal{Q}f(n)=\lambda n\left(f(n+1)-f(n)\right)+\sum_{i=1}^{n-1}\sum_{|\eta|=i}\tilde{q}_{\xi_{0}^{l}\eta}^{n}\left(f(|\eta|)-f(n)\right).$
It is a general fact about Markov processes that for any $f$ in $C_{b}(\mathbb{Z}_{+})$ that
$M(t):=f(L_{t})-f(L_{0})-\int_{0}^{t}\mathcal{Q}f(L_{s})ds$ (7.8)
is a martingale with quadratic variation
$\left\langle M\right\rangle_{t}=\int_{0}^{t}\mathcal{Q}(f^{2})(L_{s})-2f(L_{s})\mathcal{Q}f(L_{s})ds;$ (7.9)
see KL (98)[Lemma 5.1] or EK (09)[Proposition 4.1.7]. A truncation argument will enable us to take $f$ to be the identity function $Id$ (i.e. when $Id(n)=n$ for all $n\in\mathbb{Z}_{+}$) to obtain
[…]
where $M^{Id}$ corresponds to the suitable martingale defined in (7.8) with $f=Id$. We calculate directly that
$\mathcal{Q}(Id)(n)$ $=\lambda n+\sum_{i=1}^{n-1}\sum_{|\eta|=i}\tilde{q}_{\xi_{0}^{n}\eta}^{n}(i-n)\leq\lambda n,$ (7.11)
$\mathcal{Q}(Id^{2})(n)$ $=\lambda n(2n+1)+\sum_{i=1}^{n-1}\sum_{|\eta|=i}\tilde{q}_{\xi_{0}^{n}\eta}^{n}(i^{2}-n^{2}),\text{ and }$ (7.12)
$\mathcal{Q}(Id^{2})(n)-2n\mathcal{Q}(Id)(n)$ $=n\left(\lambda-\left|\tilde{q}_{\xi_{0}^{n}\xi_{0}^{n}}^{n}\right|\right)+\sum_{i=1}^{n-1}\sum_{|\eta|=i}i\tilde{q}_{\xi_{0}^{n}\eta}^{n}\leq n\lambda.$ (7.13)
Combining (7.10), (7.11), and (7.9) yields
$L_{t}$ $\leq L_{0}+\lambda\int_{0}^{t}L_{s}ds+M^{Id}(t),\text{ and }$ (7.14)
$\left\langle M^{Id}\right\rangle_{t}$ $\leq\lambda\int_{0}^{t}L_{s}ds.$ (7.15)
This allows us to compare $L$ with the solution $X$ to
$dX_{t}=\lambda X_{t-}+dM^{Id}_{t},\quad X_{0}=L_{0},$
using the observation that
$\mathbb{P}\left(\sup_{t\in[0,T]}L_{t}\geq K\right)\leq\mathbb{P}\left(\sup_{t\in[0,T]}X_{t}\geq K\right).$ (7.16)
Observe that $X_{t}=e^{\lambda t}\left(L_{0}+N_{t}\right)$ where $N_{t}=\int_{0}^{t}e^{-\lambda s}dM^{Id}(s)$. Therefore
$\sup_{t\in[0,T]}X_{t}\leq e^{\lambda T}\left(L_{0}+\sup_{t\in[0,T]}\int_{0}^{t}e^{-\lambda s}dM^{Id}(s)\right).$ (7.17)
By (7.9), (7.16) and Doob’s inequality applied to (7.17) we get that there exists $M(\varepsilon,T)$ such that
$\mathbb{P}\left(\sup_{0\leq t\leq T}l(t)\leq M(\varepsilon,T)\right)\leq\varepsilon.$
Finally, (7.7) can be obtained by the same argument above, when $L_{t}$ and $\int_{0}^{t}\mathcal{Q}f(L_{s})ds$ are replaced, respectively, by $l^{N}(k)$ and $\sum_{i=0}^{k-1}\mathcal{Q}^{N}f(l^{N}(i))$ for $\mathcal{Q}^{N}$ the generator of $l^{N}$. Indeed, the uniform control in $l$ given by Lemma 7.6, which depends on Assumptions 3.3 and 3.4, is enough for the argument above to follow as Equation 7.11 holds with the limiting transition rates $q^{n}_{\xi\eta}$ replaced by the finite $N$ transition rates $q^{N,n}_{\xi\eta}$, which follows by Assumption 3.6 (here $\xi,\eta$ are simply placeholders for the true elements used in the above equation). (7.7) implies the convergence of $\bar{G}^{N}$ in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to $G_{\tilde{Q},\lambda}^{n}$ by EK (09)[Corollary 8.9, Chapter 4] and (7.5).
The discrete-time ancestral graph $G^{N,n}=(G^{N,n}(k))_{k\in\mathbb{Z}_{+}}$ can be viewed graphically as having $n$ initial nodes (the nodes at the bottom of Figure 6). These $n$ nodes correspond to the $n$ individuals $\{\hat{X}_{i}^{N}(0)\}_{1\leq i\leq n}$ from whom we sampled the $n$ lineages at time-step 0. Viewed as a $\Theta$-valued process, this graph has the following key advantages: it get rid of the labels of all individuals, while keeping track of the set of all possible ancestral individuals (of the sample) and their parental relationships.
As mentioned, the ancestral lines $\{X_{i}^{N}\}$ defined in (3.1) is a system of coalescing random walks on the pedigree. We shall construct a system of coalescing random walks, view as $\mathbb{Z}_{+}$-valued processes, on the reduced object $G^{N,n}$ as follows.
Define $x_{i}^{N}(0):=i$ for all $1\leq i\leq n$. Suppose $x_{i}^{N}(j)$ is defined for all $j\leq k$ for induction. If $\hat{X}_{i}^{N}(k)$ experiences an outcrossing in the $k$th time-step, then $x_{i}^{N}(k+1)$ is equal to $l^{N}(k+1)$ with probability $\frac{1}{2}$ or else remains equal to $x_{i}^{N}(k)$ with probability $\frac{1}{2}$. (In the event that there is more than one outcrossing in a single time-step, it may be that each outcrossing individual coalesces in the position $l+1$; in the limited outcrossing regime, this possibility is negligible, so we do not need to consider this case in any detail.) We therefore have $x_{i}^{N}=\left(x_{i}^{N}(k)\right)_{k\in\mathbb{Z}_{+}}$ defined for all time. If any two walks $x_{i}^{N}$ and $x_{j}^{N}$ coalesce at any point in time, they are then required to remain identical for all time thereafter. This gives a family of $n$ simple random walks on $G^{N,n}$ satisfying
  1. 1.
    $x_{i}^{N}(k)\leq l^{N}(k)$ for all $k\in\mathbb{Z}_{+}$,
  2. 2.
    at any outcrossing event, each random walk follows each of the two paths available with equal probability,
  3. 3.
    and once two random walks coalesce they remain together for all time thereafter.
Lemma 7.8 is readily extended to the joint convergence of the graph with the walks thereon, namely the following lemma.
Definition 7.9.
We may also define an associated coalescent process based on these $x_{i}^{N}$. Define $\mathcal{L}^{N,n}_{G}$ to be the quenched law of the sped-up coalescent process $\Pi_{G}^{N,n}=\left(\Pi_{G}^{N,n}(t)\right)_{t\in\mathbb{R}_{+}}$ defined by
$i\sim j\text{ with respect to }\Pi_{G}^{N,n}(t)\text{ if }x_{i}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)=x_{j}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right).$
That is
$\mathcal{L}^{N,n}_{G}:=\mathbb{P}\left(\Pi^{N,n}_{G}\in\cdot\,\mid\,G^{N,n}\right).$ (7.18)
Remark 7.10.
Note that $\mathcal{L}^{N,n}_{G}$ defined in (7.18) is equal to $\mathbb{P}\left(\chi^{N,n}(\lfloor tc_{N}^{-1}\rfloor)\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ which admits no random time change, unlike that of $\mathcal{L}^{N,n}$. This is because, unlike the random walks on the pedigree, we assume being in the same node in the ancestral graph $G^{N,n}$ means instantaneous coalescence of the particles. The comparison between $\mathcal{L}^{N,n}_{G}$ and $\mathcal{L}^{N,n}$ which squares our convergence results with this discrepancy is Lemma 7.14.
We will show in Lemma 7.14 that $\mathcal{L}^{N,n}$ and $\mathcal{L}^{N,n}_{G}$ have the same weak limit. A suitable notion of continuity for random walks on $\Theta$-valued processes, established in Section 7.3, will then give the claimed result.
7.3 Random walks on $\Theta$-valued processes
To each element $g=\left(g(t)\right)_{t\in\mathbb{R}_{+}}=\left(l(t),m(t),\xi(t)\right)_{t\in\mathbb{R}_{+}}$ in the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$, we will define an associated coalescent process $\chi(g)$. Let $g(t-)=\lim_{s\uparrow t}g(s)=(l(t-),m(t-),\xi(t-))$ denote the left limit of $g$ at $t$. We suppose that $g(0)=(l_{0},m_{0},\xi_{0})$ and define, for all $1\leq i\leq l_{0}$ that $x_{i}(0)=i$. The $x_{i}$ may be extended to all time as follows:
  • On any interval $[s,t)$ on which $l$ is constant, each $x_{i}$ stays constant.
  • At each jump time $t$ of $g$, if l(t)<l(t)l(t)<l(t-), then we set
    $x_{i}(t)=\inf\{[x_{i}(t-)]_{\xi(t)}\},$
    where $a]_{\xi}$ denotes the block of $\xi$ containing $a$. If the infimum is taken over an empty set, then $x_{i}(t)=x_{i}(t-)$.
  • At each jump time $t$ of $g$, if l(t)>l(t)l(t)>l(t-), then we let $C(t)$ be an independent random variable taking value in $\{m(t),l(t)\}$ where $\mathbb{P}\left(C(t)=m(t)\right)=\mathbb{P}\left(C(t)=l(t)\right)=\frac{1}{2}$. Then for every $i$ in the set $I:=\{i:x_{i}(t-)=m(t)\}$, we let $x_{i}(t):=C(t)$. If $i$ is not in $I$, then $x_{i}(t)=x_{i}(t-)$.
The $\{x_{i}\}_{i=1}^{n}$ as $\mathbb{Z}_{+}$-valued processes take the form a family of coalescing random walks on $g$. We define their associated coalescent $\Pi^{n}(g)$ by
$i\sim j\text{ with respect to }\Pi^{n}(g)(t)\text{ if }x_{i}(t)=x_{j}(t).$
Remark 7.11.
Note that $\Pi^{N,n}_{G}$ in Definition 7.9 is equal to $\Pi^{n}(\bar{G}^{N})$, where $\bar{G}^{N}$ is the sped-up ancestral graph of Lemma 7.8.
Lemma 7.12.
Let $\mathcal{E}:=\bigsqcup_{n\in\mathbb{N}}\mathcal{E}_{n}$. The measure-valued map $\Phi:\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)\to\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right)\right)$ defined by
$\Phi(g):=\mathbb{P}\left(\Pi(g)\in\cdot\,\mid\,g\right).$
is continuous.
Proof.
Suppose that $g_{N}$ is a sequence of elements in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ that converge therein to an element $g$. We need to show that $\Phi(g_{N})$ converges weakly in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right)\right)$ to $\Phi(g)$. That is, we need to show that for any dense subset $\mathcal{F}$ of functions $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}\right))$ that $\int fd\Phi(g_{N})$ converges to $\int fd\Phi(g)$. As $\Theta$ is a locally compact Polish space, by Theorem 6.5 it suffices to check that
$\left(\Phi(g_{N})\left(\int_{0}^{\infty}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\left(\Phi(g)\left(\int_{0}^{\infty}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}$
for any collection $\left(\varphi_{i}\right)_{1\leq i\leq k}$ of elements in $C_{c}\left(\mathbb{R}_{+}\times\Theta\right)$.
As $g_{N}$ converges to $g$ and $\Theta$ is a discrete space we have that, for large enough $N$, that $g_{N}(0)$ is eventually constant. In particular, we may restrict to the case that $g_{N}(0)=(n,m_{0},\xi_{0})$ for all $N$. This allows us to consider continuity of the restriction of $\Phi$ as a map into $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ without loss of generality.
As each $\varphi_{i}$ is compactly supported on $\mathbb{R}_{+}\times\Theta$, there is a $T>0$ such that $\cup_{i}supp(\varphi_{i})\subset[0,T]\times E$. Notice that, for any $N$ that
$\left(\Phi(g_{N})\left(\int_{0}^{\infty}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}=\left(\mathbb{E}\left[\int_{0}^{T}\varphi_{i}(s,\Pi^{n}(g_{N})(s))ds\in\cdot\,\mid\,g_{N}\right]\right)_{1\leq i\leq k}.$ (7.19)
The quantity on the right of (7.19) depends only on discrete graph structure of $g_{N}$, i.e. the finite sequence of values that $g_{N}$ takes, and the edge lengths of the graph $g_{N}$, i.e. the jump times. As $\Theta$ is discrete and $g_{N}\to g$ in the Skorokhod $J_{1}$ topology, for large enough $N$ the discrete graph structure of $g_{N}$ is eventually constant. Further, the edge lengths of $g_{N}$ converge uniformly on $[0,T]$ to those of $g$. This yields the desired convergence.
As such, the coalescing random walks on $g_{N}$ and $g$ may be coupled in such a way that, on $[0,T]$, they make the same jumps at each fragmentation event as each other. In particular, we can take it so that $\Pi^{n}(g_{N})$ converges pointwise almost surely on $[0,T]$ to $\Pi^{n}(g)$. Because the integrands are bounded and convergence is pointwise almost surely on $[0,T]$, the dominated convergence theorem yields the claim.
As a corollary, we have that if $G_{N}$ is a sequence of $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$-valued random variables converging weakly to $G$, then $\Phi(G_{N})$ converges weakly in distribution to $\Phi(G)$.
7.4 Proof of Theorem 4.7
To prove Theorem 4.7, we establish a lemma demonstrating the equivalence of the quenched coalescent for the $Q$-$\lambda$ graph, as in Definition 7.3, and the quenched coalescent for an EFC, as in Definition 4.6. Before the formal statement of the lemma, we give an intuitive explanation for why the two constructions agree. Fix any finite sample of at most $n$ particles and follow their coalescing random-walk trajectories. Under both the $Q$-$\lambda$ graph and the EFC, coalescences of the walks occur with the same structures and rates and fragmentations of each walk occur with the same rate. Because these rates uniquely determine the law of the induced coalescent on partitions and because both the $Q$-$\lambda$ graph and EFC give the same amount of information on the particle trajectories, the quenched coalescent laws coincide. The lemma that follows makes this argument rigorous via a joint coupling of the $Q$-$\lambda$ graph, the EFC, and random walks thereon.
Lemma 7.13.
Suppose that $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is a consistent family of generators for a $\Xi$-coalescent for some $\Xi$ in $\mathcal{M}_{1}\left(\Delta\right)$. Let $\Xi^{\prime}=\Xi-\delta_{\mathbf{0}}\Xi\left(\mathbf{0}\right)$. Let $\Pi$ be an EFC process with characteristics $c_{k}=\Xi(\mathbf{0})$, $\nu_{Coag}=\nu_{\Xi^{\prime}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$. Then the random measures $\mathcal{L}_{\Pi}^{n}$ and $\mathcal{L}_{Q,\lambda}^{n}$ are equal in distribution.
Proof.
For a given $Q$-$\lambda$ graph $G=\left((l(t),m(t),\xi(t))\right)_{t\geq 0}$ and an EFC process $\Pi=\left(\Pi(t)\right)_{t\geq 0}$, we let $x^{G}=\left(x_{i}^{G}\right)_{1\leq i\leq n}$ and $x^{\Pi}=\left(x_{i}^{\Pi}\right)_{1\leq i\leq n}$ denote the family of coalescing random walks on $G$ and $\Pi$, respectively. Let $\chi^{n}_{G}$ and $\chi^{n}_{\Pi}$ denote the associated coalescents for $x^{G}$ and $x^{\Pi}$, respectively. We describe a coupling of $\Pi$, $G$, $x^{\Pi}$, and $x^{G}$ together so that the claim holds. Specifically, it will be such that $\chi^{n}_{G}=\chi^{n}_{\Pi}$ pointwise, and such that $\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,\Pi\right)=\mathbb{P}\left(\chi^{n}_{\Pi}\in\cdot\,\mid\,G\right)$, which together prove the claim. The construction follows that for the EFC (Ber, 04, Section 3.2).
We take $\Pi(0)=\xi_{0}$, $G(0)=\left(n,0,\xi_{0}^{n}\right)$, $x^{G}(0)=x^{\Pi}(0)=\left(1,2,\ldots,n\right)$, and $\xi(0)=\xi_{0}^{n}$. Let $P_{C}=\left((t,\xi^{(C)}(t))\right)_{t\geq 0}$ and $P_{F}=\left(\left(t,\xi^{(F)}(t),k(t)\right)\right)_{t\geq 0}$ be two independent Poisson point processes (PPPs) on the same filtration. The atoms of $P_{C}$ are points in $\mathbb{R}_{+}\times\mathcal{E}_{\infty}$ and $P_{C}$ has intensity measure $dt\otimes C$. The atoms of $P_{F}$ are points in $\mathbb{R}_{+}\times\mathcal{E}_{\infty}\times\mathbb{N}$ and has intensity measure $dt\otimes F\otimes\#$, where #\# is the counting measure on $\mathbb{N}$. We utilize $P_{F}$ and $P_{C}$ to construct $(\Pi,G,x^{G},x^{\Pi})$ jointly.
To this end, we establish notation for the restriction of partitions of $\mathbb{N}$ to finite subsets. For $\xi\in\mathcal{E}_{\infty}$ and a finite subset $I=\left\{i_{j}\right\}_{1\leq j\leq l}$ of $\mathbb{N}$, we define $\xi(I)$ to be the partition of $[l]$ defined by
$j\sim_{\xi(I)}k\text{ if }i_{j}\sim_{\xi}i_{k}.$
With this notation, the coupling may be described as follows:
  1. 1.
    If $t$ is neither an atom time of $P_{F}$ nor $P_{C}$, then $\Pi(t)=\Pi(t-)$, G(t)=G(t)G(t)=G(t-), $x^{\Pi}(t)=x^{\Pi}(t-)$, and $x^{G}(t)=x^{G}(t-)$.
  2. 2.
    If $t$ is an atom time of $P_{C}$, then
    1. (a)
      $\Pi(t)=Coag(\Pi(t-),\xi^{(C)}(t))$,
    2. (b)
      $G(t)=\left(\left|\xi^{(C)}(t)\left(\{x_{i}^{\Pi}(t-)\}_{1\leq i\leq n}\right)\right|\right),m(t-),\xi^{(C)}(t)\left(\{x_{i}^{\Pi}(t-)\}_{1\leq i\leq n}\right)$,
    3. (c)
      for all $1\leq i\leq n$, $x_{i}^{\Pi}(t)=\inf\left[x_{i}^{\Pi}(t-)\right]_{\xi^{(C)}(t)}$, where $a]_{\xi}$ denotes the block in $\xi$ to which $a$ belongs,
    4. (d)
      and $x_{i}^{G}(t)$ is equal to the index of the block of $\xi^{(C)}(t)\left(\{x_{i}^{\Pi}(t-)\}_{1\leq i\leq n}\right)$ to which $x_{i}^{\Pi}(t-)$ belongs, when ordered by their least element.
  3. 3.
    If $t$ is an atom time of $P_{F}$, then
    1. (a)
      $\Pi(t)=Frag(\Pi(t-),\xi^{(F)}(t),k(t))$,
    2. (b)
      if $k(t)\notin\{x_{i}^{\Pi}(t)\}_{1\leq i\leq n}$, then G(t)=G(t)G(t)=G(t-), $x^{\Pi}(t)=x^{\Pi}(t-)$, and $x^{G}(t)=x^{G}(t-)$,
    3. (c)
      and if $k(t)\in\{x_{i}^{\Pi}(t)\}_{1\leq i\leq n}$, then $G(t)=\left(l(t-),m(t),\xi(t-)\right)$, where $m(t)$ is equal to $x_{i}^{G}(t-)$ (for any $i$ such that $x_{i}^{\Pi}(t-)=k(t)$; check that this is indeed well-defined); for $H$ an independent Bernoulli 010-1 random random variable with $\mathbb{P}\left(H=0\right)=\mathbb{P}\left(H=1\right)=\frac{1}{2}$ and if the only non-singleton block of $\xi^{(F)}(t)$ is $\{r,s\}$, then
      $x_{i}^{\Pi}(t)=\begin{cases}x_{i}^{\Pi}(t-)&,\text{ if }k(t)\neq x_{i}^{\Pi}(t-)\text{ or }H=0\\ s&,\text{ if }k(t)=x_{i}^{\Pi}(t-),H=1\end{cases}$
      and
      $x_{i}^{G}(t)=\begin{cases}x_{i}^{G}(t-)&,\text{ if }x_{i}^{\Pi}(t)=x_{i}^{\Pi}(t-)\\ l(t)&,\text{ if }x_{i}^{\Pi}(t)=s\end{cases}.$
By construction of $P_{F}$ and $P_{C}$ it is manifest that $\Pi$ is an EFC as described and $G$ is a $Q$-$\lambda$ graph. Further, it is indeed the case that $\chi^{n}_{G}=\chi^{n}_{\Pi}$ pointwise. As the structure of coalescent events and fragmentation events for the $x^{\Pi}$ are fully determined by $G$ we indeed have that $\mathbb{P}\left(\chi^{n}_{\Pi}\,\mid\,\Pi\right)=\mathbb{P}\left(\chi^{n}_{\Pi}\,\mid\,G\right)$. This yields the claim.
We now show that if $\mathcal{L}^{N,n}_{G}$ converges in distribution as a random measure to a random measure $\mathcal{L}$, then so does $\mathcal{L}^{N,n}$. In particular, these two sequences have the same distributional limit. The spirit of the argument is really that the path of elements in $\mathcal{E}_{n}$ for $\bar{\chi}^{N,n}$ and $\Pi^{N,n}_{G}$ are the same with high probability, and that the two differ on a set of negligible measure as $N\to\infty$. As we compare these measures via integral functionals of compactly supported test functions, in the sense of Theorem 6.5, this is sufficient to show that the differences in their evaluations of these integral functions vanishes as $N$ tends to infinity.
Lemma 7.14.
Suppose that, as $N\to\infty$, that $c_{N}\to 0$, $d_{N}c_{N}^{-1}\to\lambda<\infty$, and $\mathcal{L}^{N,n}_{G}$ converges in distribution in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$ to $\mathcal{L}$. Then $\mathcal{L}^{N,n}$ also converges in distribution to $\mathcal{L}$ in $\mathcal{M}_{1}\left(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)\right)$.
Proof.
We wish to apply Theorem 6.5 to demonstrate the claim. To this end, fix $\{\varphi_{i}\}_{i=1}^{k}\subset C_{c}(\mathbb{R}_{+}\times\mathcal{E}_{n})$ a finite collection of test functions. By Theorem 6.5 and the assumption that $\mathcal{L}^{N,n}_{G}\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd}}}}\,\mathcal{L}$ we have that
$\left(\mathcal{L}^{N,n}_{G}\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}$ $=\left(\mathbb{E}\left[\int_{\mathbb{R}_{+}}\varphi_{i}\left(s,\Pi^{N,n}_{G}(s)\right)ds\,\mid\,G^{N}\right]\right)_{1\leq i\leq k}$ (7.20)
$\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\left(\mathcal{L}\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)\right)_{1\leq i\leq k}.$ (7.21)
For the claim to hold, by Theorem 6.5 and Equation (7.20), it suffices to show that, for any one of the $\varphi_{i}$, that
$\mathcal{L}^{N,n}\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)-\mathcal{L}^{N,n}_{G}\left(\int_{\mathbb{R}_{+}}\varphi_{i}(s,x(s))ds\right)$ (7.22)
converges in probability to 0. To this end, we fix $\varphi$ in $C_{c}(\mathbb{R}_{+}\times\mathcal{E}_{n})$ and show precisely this convergence. We proceed by analyzing how $\bar{\chi}^{N,n}$ and $\Pi^{N,n}_{G}$ differ.
By construction of $\Pi_{G}^{N,n}$ we have that, if $x_{i}^{N}(t)=x_{j}^{N}(t)$, then $i$ and $j$ belong to the same block in $\Pi^{N,n}_{G}(t)$. However, for the corresponding random walks on the pedigree $\hat{X}_{i}^{N}(\lfloor tc_{N}^{-1}\rfloor)=\hat{X}_{j}^{N}(\lfloor tc_{N}^{-1}\rfloor)$ does not imply that $i$ and $j$ belong to the same blocks in $\bar{\chi}^{N,n}(t)$. Instead, it means that $\chi^{N,n}(\lfloor tc_{N}^{-1}\rfloor)\notin\mathcal{E}_{n}$, and so we are experiencing the random time-change where $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$. The two blocks containing $i$ and $j$ may either coalesce before $\chi^{N,n}$ reenters $\mathcal{E}_{n}$ or else disperse. We claim that dispersal occurs with probability tending to $1$ as $N\to\infty$, and that the time $T$ until either coalescence or dispersal satisfies $c_{N}T$ converges in probability to 0. This will demonstrate further, at all times where $S(\lfloor tc_{N}^{-1}\rfloor)=\lfloor tc_{N}^{-1}\rfloor$, that $\Pi_{G}^{N,n}(t)=\bar{\chi}^{N,n}(t)$ with probability tending to $1$ as $N\to\infty$, for each $t\in(0,\infty)$.
Indeed, by finiteness of the partition sizes it suffices to show this with the partition $\{(\{1\},\{2\})\}$. Observe that the number $U$ of selfing events in the ancestral line of the sample before the ancestral lines the sample experience an outcrossing satisfies
$\mathbb{P}\left(U=r\right)=\alpha_{N}^{r}(1-\alpha_{N}).$
At each selfing event, the two sample lineages have probability $\frac{1}{2}$ of coalescing. Therefore, the probability of not coalescing before the outcrossing is
$\sum_{r=0}^{\infty}\alpha_{N}^{r}(1-\alpha_{N})2^{-r}=\frac{1-\alpha_{N}}{1-\frac{\alpha_{N}}{2}}.$ (7.23)
By Assumption 3.4 we know that $\alpha_{N}\to 1$, which shows that the non-coalescence probability of Equation (7.23) converges to 0 as $N\to\infty$. That $c_{N}T$ converges to 0 as $N\to\infty$ follows from Assumption 3.3 and the fact that the number of time-steps until a coalesce event or a dispersal is geometric with success probability $\frac{K_{N}}{N}\left(\frac{\alpha_{N}}{2}+1-\alpha_{N}\right)$. By a direct computation this is $O(c_{N})$ in expectation. The $\frac{K_{N}}{N}$ term comes from the probability that the individual containing the two sample lineages is a child, $\frac{\alpha_{N}}{2}$ is the probability that a selfing event occurs and yields a coalescence, and $1-\alpha_{N}$ is the probability of seeing an outcrossing.
We have established thus far that $\Pi_{G}^{N,n}(t)=\bar{\chi}^{N,n}(t)$ for all $t$ in which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ with probability 1o(1)1-o(1). Further, we have shown that the measure of time $t$ on which $S(\lfloor tc_{N}^{-1}\rfloor)\neq\lfloor tc_{N}^{-1}\rfloor$ converges to 0 in distribution as $N$ goes to infinity. Therefore we have convergence in probability to 0 of Equation (7.22), as needed. ∎
We now proceed with the proof of Theorem 4.7.
Proof.
Notice by Lemma 7.13, it suffices to show convergence as random measures of $\mathcal{L}^{N,n}$ to $\mathcal{L}_{Q,\lambda}^{n}$. By Lemma 7.8 we have that $G^{N,n}$ converges weakly in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to a $Q$-$\lambda$ graph $G$. By Lemma 7.12 we know that the map $\Phi$ is continuous. Continuity preserves weak convergence and so $\mathcal{L}_{G}^{N,n}=\Phi(G^{N})$ converges weakly as a random measure to $\mathcal{L}_{Q,\lambda}^{n}=\Phi(G)$. By Lemma 7.14 we therefore have that $\mathcal{L}^{N,n}$ converges weakly as a random measure to $\mathcal{L}_{Q,\lambda}^{n}$. ∎
Acknowledgements
Research supported by National Science Foundation grants DMS-2534011, DMS-2532674, DMS-2348164 and DMS-2152103. We appreciate Adam Jakubowski for pointing out references BK10a ; BK10b and the lack of method to construct a dense family of continuous bounded functions on the (non-locally compact) Skorokhod space.
Appendix Appendix A Convergence Criteria for the diploid model with selfing and overlapping generations
To provide convergence criteria for our diploid exchangeable model with selfing and overlapping generations, we proceed along the lines of [37] and [10, 1]. We define the offspring count $V_{i}$ for the $i$th individual to count the number of genetic descendents among the $2N$ genes in the next time-step that can trace their lineage back to the $i$th individual. Formally, we have
$V_{i}=\sum_{j}V_{i,j}+2V_{i,i}.$
Selfing counts twice as each individual contains two sample lineages. Denote by
$V_{(1)}\geq V_{(2)}\geq\ldots\geq V_{(N)}$
the ranked version of the total offspring numbers $\left(V_{i}\right)_{1\leq i\leq N}$ and by
$\Phi_{N}:=\mathscr{L}\left(\frac{V_{(1)}}{2N},\frac{V_{(2)}}{2N},\ldots,\frac{V_{(N)}}{2N},0,0,\ldots\right)$
the law of the ranked offspring frequencies, viewed as a probability measure on the infinite ordered simplex $\Delta$.
Lemma Appendix A.1.
Suppose, as $N\to\infty$, that Assumption 3.3 holds and that
$\frac{1}{2c_{N}}\Phi_{N}(dx)\to\frac{1}{\left\langle x,x\right\rangle}\Xi^{\prime}(dx)$ (Appendix A.1)
vaguely on $\Delta\setminus\{\mathbf{0}\}$, where $\Xi^{\prime}$ is a probability measure on $\Delta$. Then Assumption 3.6 holds where, for each $n\geq 2$, $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
Proof.
Let $\mathit{P}_{N,n}=\left(\mathit{p}_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$ denote the one-step transition matrix of the ancestral process after applying the haploid map $F_{\rm hap}$ of Equation (3.10) so that
$\mathit{p}_{\xi\eta}^{N,n}=\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\,\mid\,\chi^{N,n}(0)=\xi\right).$
To prove the claim we need to show that $c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)$ converges to a matrix $Q_{n}$ that is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$.
Without loss of generality, consider $\xi=\xi_{0}^{b}$ for some $1\leq b\leq n$, and let $\eta$ be of $(k_{1},\ldots,k_{r};s)$-type. That is, $\eta$ is obtained from $\xi_{0}^{b}$ by keeping $s$ of the blocks as singletons and coalescing the remaining bsb-s blocks into $r$ blocks of sizes $k_{1},\ldots,k_{r}$. Fix $\varepsilon>0$ and define $A_{\varepsilon}:=\{x\in\Delta:\langle x,x\rangle>\varepsilon\}$, with complement $A_{\varepsilon}^{\mathrm{c}}:=\Delta\setminus A_{\varepsilon}$. We decompose
$Q_{N,n}=c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)=\Phi_{N}(A_{\varepsilon})\,Q_{N,n}^{\varepsilon}+\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\,Q_{N,n}^{\varepsilon,\mathrm{c}},$ (Appendix A.2)
where $Q_{N,n}^{\varepsilon}$ and $Q_{N,n}^{\varepsilon,\mathrm{c}}$ are the conditional transition matrices given $x\in A_{\varepsilon}$ and $x\in A_{\varepsilon}^{\mathrm{c}}$, respectively. By Assumption (Appendix A.1),
$\frac{\Phi_{N}(A_{\varepsilon})}{2c_{N}}\to\int_{A_{\varepsilon}}\frac{1}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
and hence $\Phi_{N}(A_{\varepsilon})\to 0$ and $\Phi_{N}(A_{\varepsilon}^{\mathrm{c}})\to 1$ as $N\to\infty$ by Assumption 3.3.
For $x\in A_{\varepsilon}$, denote by
$\Phi_{b;(k_{1},\ldots,k_{r};s)}(x):=\sum_{i_{1},\ldots,i_{r}\ \mathrm{distinct}}\Big(\prod_{j=1}^{r}x_{i_{j}}^{k_{j}}\Big)\,T^{(r)}_{r,s}(x_{i_{1}},\ldots,x_{i_{r}}),$
the probability that $b$ blocks thrown uniformly on $[0,1]$ with a paintbox $x$ form a partition of type $k_{1},\ldots,k_{r};s$. Here $T^{(r)}_{r,s}$ is the inclusion–exclusion term enforcing that the $s$ singletons fall into distinct intervals from [10, p. 41, Appendix A]. Then, conditional on $x$, we have
$\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)=\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)+\Delta_{N}(x),\qquad\text{with }|\Delta_{N}(x)|\leq\frac{K_{b}}{N}.$
That $|\Delta_{N}(x)|$ is $O(\frac{1}{N})$ follows simply from the observation that the probability of the $n\ll N\sum_{i}x_{i}$ sample lineages falling into each of the intervals of $x$ are asymptotically independent, with corrections at worst of order $\frac{1}{N}$, which can be uniformly controlled by compactness of $\Delta$. Therefore,
$\frac{1}{c_{N}}\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\mathbb{P}\left(F_{\rm hap}(\chi^{N,n}(1))=\eta\mid x,\chi^{N,n}(0)=\xi_{0}^{b}\right)\right]=\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)\right]+o(1),$
since the $\frac{1}{N}$ coupling error is multiplied by $\Phi_{N}(A_{\varepsilon})/c_{N}=O_{\varepsilon}(1)$. Here $O_{\varepsilon}(1)$ simply means that, the term is $O(1)$ for any fixed $\varepsilon>0$. By Assumption (Appendix A.1), vague convergence on $\Delta\setminus\{\mathbf{0}\}$ yields
$\lim_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\left[\mathbf{1}_{A_{\varepsilon}}\,\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)\right]=2\int_{A_{\varepsilon}}\frac{\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
and letting $\varepsilon\downarrow 0$ by monotone convergence gives
$2\int_{\Delta\setminus\{\mathbf{0}\}}\frac{\Phi_{b;(k_{1},\ldots,k_{r};s)}(x)}{\langle x,x\rangle}\,\Xi^{\prime}(dx),$
which coincides with the transition for $\Xi\!\restriction_{\Delta\setminus\{\mathbf{0}\}}=2\Xi^{\prime}$.
For $x\in A_{\varepsilon}^{\mathrm{c}}$ we have $\langle x,x\rangle\leq\varepsilon$, hence $\max_{i}x_{i}\leq\|x\|_{2}\leq\varepsilon^{1/2}$. Then
$\sum_{i}x_{i}^{3}\;\leq\;(\max_{i}x_{i})\sum_{i}x_{i}^{2}\;\leq\;\varepsilon^{1/2}\sum_{i}x_{i}^{2},\qquad\Big(\sum_{i}x_{i}^{2}\Big)^{2}\;\leq\;\varepsilon\,\sum_{i}x_{i}^{2}.$
Consequently,
$\mathbb{P}(\text{non-binary}\mid x)\;\leq\;C_{b}\!\left(\sum_{i}x_{i}^{3}+\Big(\sum_{i}x_{i}^{2}\Big)^{2}\right)\;\leq\;C_{b}\big(\varepsilon^{1/2}+\varepsilon\big)\sum_{i}x_{i}^{2}\;\leq\;2C_{b}\,\varepsilon^{1/2}\sum_{i}x_{i}^{2}$
for some constant $C_{b}$ depending only on $b$. Therefore, using $\lim_{N\to\infty}\frac{1}{c_{N}}\mathbb{E}\!\big[\sum_{i}x_{i}^{2}\big]=2\Xi^{\prime}(\Delta\setminus{\{\mathbf{0}}\})$,
$\limsup_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(\text{non-binary}\mid x)\big]\;\leq\;2C_{b}\,\varepsilon^{1/2}\,\limsup_{N\to\infty}\frac{1}{c_{N}}\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]\;=\;4C_{b}\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})\,\varepsilon^{1/2}.$
In particular,
$\lim_{\varepsilon\downarrow 0}\;\limsup_{N\to\infty}\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(\text{non-binary}\mid x)\big]=0.$
For any unordered pair $\{a,b\}$,
$\mathbb{P}(\{a,b\}\text{ coalesce}\mid x)=\sum_{i}x_{i}^{2}+O\!\Big(\sum_{i}x_{i}^{3}\Big),$
and on $A_{\varepsilon}^{\mathrm{c}}$ the error satisfies
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{3}\big]\;\leq\;\varepsilon^{1/2}\,\frac{1}{c_{N}}\,\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]\;\xrightarrow[N\to\infty]{}\;2\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})\,\varepsilon^{1/2}.$
Thus
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\mathbb{P}(\{a,b\}\text{ coalesce}\mid x)\big]=\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]+o_{\varepsilon}(1)$
with $o_{\varepsilon}(1)\to 0$ as $\varepsilon\downarrow 0$ uniformly in $N$. Decomposing,
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]=\frac{1}{c_{N}}\,\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]-\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}}\sum_{i}x_{i}^{2}\big].$
By the normalization and the vague convergence assumption (Appendix A.1),
$\frac{1}{c_{N}}\,\mathbb{E}\!\Big[\sum_{i}x_{i}^{2}\Big]\;\to\;2\,\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\}),\qquad\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}}\sum_{i}x_{i}^{2}\big]\;\to\;2\,\Xi^{\prime}(A_{\varepsilon}),$
hence
$\frac{1}{c_{N}}\,\mathbb{E}\!\big[\mathbf{1}_{A_{\varepsilon}^{\mathrm{c}}}\sum_{i}x_{i}^{2}\big]\;\to\;2\,\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})-2\,\Xi^{\prime}(A_{\varepsilon})\xrightarrow[\varepsilon\downarrow 0]{}\;0.$
By the normalization so that all binary mergers occur at rate $2c_{N}$, the remaining $2-2\Xi^{\prime}(\Delta\setminus\{\mathbf{0}\})$ mass is concentrated on binary coalescences, which gives the Kingman part of the $\Xi$-coalescent.
Combining the contributions from $A_{\varepsilon}$ and $A_{\varepsilon}^{\mathrm{c}}$ and setting $q_{n}(\xi,\xi)=-\sum_{\eta\neq\xi}q_{n}(\xi,\eta)$, we conclude that
$c_{N}^{-1}\left(\mathit{P}_{N,n}-I\right)\ \to\ Q_{n},$
where $Q_{n}$ is the generator of an $n$-$\Xi$-coalescent with $\Xi=2\Xi^{\prime}$. This completes the proof. ∎
Remark Appendix A.2.
[10, Theorem 1.1] establishes that, if Assumption 3.3 and (Appendix A.1) hold, then the rescaled ancestral process converges to the $\Xi$-coalescent with $\Xi=2\Xi^{\prime}\circ\varphi^{-1}$, where $\varphi$ denotes the halving map $\varphi:\Delta\to\Delta$ defined by
$\left(x_{1},x_{2},\ldots\right)\mapsto\left(\frac{1}{2}x_{1},\frac{1}{2}x_{1},\frac{1}{2}x_{2},\frac{1}{2}x_{2},\ldots\right).$
Note that this $\Xi$ (which is under no selfing) is different from the $\Xi=2\Xi^{\prime}$ in the annealed convergence in Theorem 3.9. This difference is expected because when $\alpha_{N}\not\to 1$, the annealed limit of the rescaled ancestral process will not have $Q_{n}$ as its generator. The haploid map $F_{\rm hap}$ in (3.10), which appears in the definition of $Q_{n}$, makes $Q_{n}$ the generator of the limiting process where two blocks that enter the same individual always coalesce instantaneously in the limit. When $\alpha_{N}\not\to 1$, blocks of the ancestral process that enter the same individual may disperse before coalescing, and this is captured in the true limiting generator.
We refer the interested reader to [10, Appendix A] for several equivalent formulations to Equation (Appendix A.1).
Appendix Appendix B Proof of the Theorem 3.9
We produce a small lemma here to show that, for our notion of weak convergence in distribution that the intensity measure map sending a random measure $\mu$ to its intensity measure $\mathbb{E}\left[\mu\right]$ is continuous.
Lemma Appendix B.1.
Suppose that $\mu_{N}$ is a sequence of random measures on $\mathcal{D}\left(\mathbb{R}_{+},E\right)$ that converge weakly in distribution, in the sense of Definition 6.1, and that $E$ is a locally compact Polish space. Then the sequence of intensity measures $\mathbb{E}\left[\mu_{N}\right]$ converge weakly in distribution to $\mathbb{E}\left[\mu\right]$ when $C_{b}\left(\mathcal{D}\left(\mathbb{R}_{+},E\right)\right)$ is given the topology of uniform convergence on compacta.
Proof.
Recall that for a random measure $\mu$ that the intensity measure $\mathbb{E}\left[\mu\right]$ is defined as the unique measure for which
$\mathbb{E}\left[\mu(f)\right]=\mathbb{E}\left[\mu\right](f)$
for any $f$ in $C_{b}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$.
The topology on $\mathcal{M}_{1}(\mathcal{D}\left(\mathbb{R}_{+},E\right))$ is such that the evaluation maps $T_{f}$ sending $\mu$ to $\mu(f)$ is continuous and bounded. By the continuous mapping theorem therefore $T_{f}(\mu_{N})=\mu_{N}(f)$ converges in distribution to $T_{f}(\mu)=\mu(f)$. Because $T_{f}$ is continuous and bounded, convergence in distribution implies convergence in expectation. This yields the claim.
With Theorem 4.7 in hand, we are ready to prove Theorem 3.9.
Proof.
Note that the assumptions of Theorem 3.9 are the same as those of Theorem 4.7. Consequently, by Theorem 4.7 we have that the sequence of random measures $\mathcal{L}^{N,n}$ converges weakly in distribution to $\mathcal{L}_{\Pi}^{n}$ where $\Pi$ is an EFC with coalescence measure $\Xi$ on $\Delta$, no non-binary fragmentation, and binary fragmentation with rate $\lambda$. By Lemma Appendix B.1, the intensity measure mapping is continuous, and so $\mathbb{E}\left[\mathcal{L}^{N,n}\right]=\mathbb{P}_{\xi_{0}^{n}}\left(\bar{\chi}^{N,n}\in\cdot\right)$ converges weakly to $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$. It suffices, therefore, to show that $\mathbb{E}\left[\mathcal{L}_{\Pi}^{n}\right]$ is the law of an $n$-$\Xi$-coalescent.
Consider a fixed sample $m\leq n$ of particles located in the EFC at time $t$. Then, regardless of the position of the $m$ particles, when we anneal over the EFC the times at which these $m$ particles coalesce is governed by $Q_{m}$, the projection of the generator $Q$ on $\mathcal{E}_{\infty}$ associated to $\Xi$ by Remark 3.8. But this is precisely the definition of the generator of an $m$-$\Xi$-coalescent in its initial state. This gives the claim. ∎
Appendix Appendix C Combinatorics for the general model
Let $N$ be fixed and consider the general diploid overlapping-generations model described at the beginning of Section 2. At time step 0, let $K:=K_{N}^{(0)}$, $S:=S_{N}^{(0)}=\sum_{i=1}^{N}V^{(0)}_{i,i}$, $v_{i}:=V^{(0)}_{i,i}$ and $u_{i}:=\sum_{j\neq i}V^{(0)}_{i,j}$ for $i\in[N]$.
Lemma Appendix C.1 (One-step triple coalescence).
The probability that three distinct sampled lineages from three distinct individuals at time 0 coalesce into a single ancestor in one step (to time $1$ in the past) is
$\mathfrak{c}_{3}:=\mathbb{P}_{\xi_{0}^{3}}\left(\chi^{N,n}(1)=\mathbf{1}_{3}\right)=\frac{1}{\binom{N}{3}}\,$ $\mathbb{E}\!\bigg[\frac{N-K}{N}\sum_{i=1}^{N}\!\left(\frac{1}{4}\binom{v_{i}}{2}+\frac{1}{8}\,v_{i}u_{i}+\frac{1}{16}\binom{u_{i}}{2}\right)$
$\;+\;\sum_{i=1}^{N}\!\left(\frac{1}{8}\binom{v_{i}}{3}+\frac{1}{16}\binom{v_{i}}{2}\,u_{i}+\frac{1}{32}\,v_{i}\binom{u_{i}}{2}+\frac{1}{64}\binom{u_{i}}{3}\right)\bigg$]. (Appendix C.1)
Here the outer expectation is over the joint law of $(K_{N}^{(0)},V^{(0)})$; $\alpha_{N}$ influences $\mathfrak{c}_{3}$ only through this law and does not appear explicitly in (Appendix C.1).
Proof.
Sample uniformly without replacement three distinct individuals at time 0; there are $\binom{N}{3}$ unordered choices. Condition first on the realized newborn set $B$ (of size $K$), the carried-over set C~C (of size NKN-K), and the parentage matrix $V^{(0)}$.
There are exactly two ways to obtain a one-step triple merger:
(A) Two newborns share a parent $i$ and the third sampled individual is that parent $i$ carried over. Fix $i\in C$. Among the $n_{i}:=v_{i}+u_{i}$ newborns that have $i$ as a parent, there are $\binom{v_{i}}{2}$ pairs where both are selfed by $i$, $v_{i}u_{i}$ mixed pairs, and $\binom{u_{i}}{2}$ pairs with both outcrossing via $i$. Tracing one step back, to coalesce with the gene copy traced from the carried-over $i$, each newborn must choose the same copy in $i$: probabilities $1/2$ (selfed) and $1/4$ (outcrossed), independently of the carried-over copy choice. Hence the Mendelian factors are $1/4$, $1/8$, and $1/16$, respectively. Summing over $i\in C$ gives the conditional contribution
$\frac{1}{\binom{N}{3}}\sum_{i\in C}\!\left(\frac{1}{4}\binom{v_{i}}{2}+\frac{1}{8}\,v_{i}u_{i}+\frac{1}{16}\binom{u_{i}}{2}\right)$.
(B) Three newborns share a common parent $i$ (not necessarily carried over). For a fixed $i\in[N]$, the three newborns can be all selfed ($\binom{v_{i}}{3}$ triples), two selfed and one outcrossed ($\binom{v_{i}}{2}u_{i}$ triples), one selfed and two outcrossed ($v_{i}\binom{u_{i}}{2}$ triples), or all outcrossed ($\binom{u_{i}}{3}$ triples). To coalesce one step back, all three lineages must choose the same copy in $i$, yielding Mendelian factors $1/8$, $1/16$, $1/32$, and $1/64$, respectively. Summing over $i$ gives the conditional contribution
$\frac{1}{\binom{N}{3}}\sum_{i=1}^{N}\!\left(\frac{1}{8}\binom{v_{i}}{3}+\frac{1}{16}\binom{v_{i}}{2}\,u_{i}+\frac{1}{32}\,v_{i}\binom{u_{i}}{2}+\frac{1}{64}\binom{u_{i}}{3}\right)$.
Adding (A) and (B) yields the conditional probability of triple coalescent given $V^{(0)},B,C$ as
$\frac{1}{\binom{N}{3}}\left[\sum_{i\in C}\!\left(\frac{1}{4}\binom{v_{i}}{2}+\frac{1}{8}\,v_{i}u_{i}+\frac{1}{16}\binom{u_{i}}{2}\right)+\sum_{i=1}^{N}\!\left(\frac{1}{8}\binom{v_{i}}{3}+\frac{1}{16}\binom{v_{i}}{2}\,u_{i}+\frac{1}{32}\,v_{i}\binom{u_{i}}{2}+\frac{1}{64}\binom{u_{i}}{3}\right)\right$].
Now average over the uniformly random carried-over set $C$ of size NKN-K, and finally, take expectation with respect to $(K_{N}^{(0)},V^{(0)})$ to obtain (Appendix C.1). ∎
By $\xi\prec\eta$ we denote that $\eta$ is obtainable from $\xi$ via a single binary merger of blocks of $\xi$. In the following lemma we calculate the one-step transition probabilities for non-binary mergers, and the one-step exit probability for a state $\xi$.
Lemma Appendix C.2.
Suppose that $\mathfrak{c}_{3}\in o(c_{N})$ as $N\to\infty$. Then the one-step transition probability $\mathit{p}_{\xi\eta}^{N,n}$ for any $\eta$ that cannot be obtained from $\xi$ via a single binary merger, i.e. where $\xi\not\prec\eta$ , satisfies $\mathit{p}_{\xi\eta}^{N,n}c_{N}^{-1}\to 0$. Further, $\mathit{p}_{\xi\xi}^{N,n}=1-2\binom{|\xi|}{2}c_{N}+o(c_{N})$.
Proof.
Note that there are $\binom{|\xi|}{2}$ possibly binary mergers for any given $\xi$, and each is equally likely by exchangeability. Therefore, it suffices to show that the exit probability after factoring through the haploid map $F_{\rm hap}$ for any $\xi$ is $2c_{N}\binom{|\xi|}{2}+o(c_{N})$. This will follow by finiteness of the state space if for any $\eta$ not obtainable by a single binary merger from $\xi$ that the one-step transition probability from $\xi$ to $\eta$ is $o(c_{N})$, which we show below. We proceed by a monotonicity argument.
Let $r$ denote the largest number of blocks of $\xi$ that are coalesced together into a single block of $\eta$. Suppose to start that $r\geq 3$ after factoring through the haploid map. For any subset of size $3$ of these $r$ blocks, of which there are $\binom{r}{3}$, the probability of that given subset coalescing in a single time-step is $\mathfrak{c}_{3}$. Therefore the one-step transition probability $\mathit{p}_{\xi\eta}^{N,n}$ that $\eta$ is obtained from $\xi$ is at most $\binom{r}{3}\mathfrak{c}_{3}$, which is $o(c_{N})$ by assumption. This shows that any $\eta$ obtained by coalescing $r\geq 3$ blocks of $\xi$ into a single block happens with $o(c_{N})$ probability.
Suppose then that $r=2$ and that $\xi\not\prec\eta$. Then there are at least four blocks $C_{1},C_{2},C_{3},C_{4}$ of $\xi$ coalesced as pairs $C_{1}\cup C_{2},C_{3}\cup C_{4}$ in $\eta$. We denote this transition by $\mathbb{P}\left(2,2\to 1,1\right)$. Note that any pairwise coalescence event, conditional on $V$, is bounded above by
$\frac{1}{4N^{2}}\sum_{i=1}^{N}(V_{i})_{2}.$ (Appendix C.2)
Therefore, where we condition on there being no triple merger, we have that
$\mathbb{P}\left(2,2\to 1,1\,\mid\,V,\text{ no triple}\right)\leq\mathbb{E}\left[\left(\frac{1}{4N^{2}}\sum_{i=1}^{N}(V_{i})_{2}\right)^{2}\right].$
By averaging over $V$ we then have
$\mathbb{P}\left(2,2\to 1,1\right)\leq\binom{4}{3}\mathfrak{c}_{3}+\mathbb{E}\left[\frac{1}{16N^{4}}\left(\sum_{i=1}^{N}(V_{i})_{2}\right)^{2}\right].$
The second summand is $O(c_{N}^{2})\subset o(c_{N})$ by a direct comparison, and by assumption $\mathfrak{c}_{3}$ is $o(c_{N})$. This shows that when $r=2$ with $\xi\not\prec\eta$ that $\mathit{p}_{\xi\eta}^{N,n}\in o(c_{N})$.
Summing over the single-pair coalescence probabilities among the $\binom{|\xi|}{2}$ unordered pairs yields the exit rate $2\binom{|\xi|}{2}c_{N}+o(c_{N})$, which completes the proof.